### v1

# A Digital twin for Diesel Engines: Operator-infused PINNs with Transfer Learning for Engine Health Monitoring

Abstract
Improving diesel engine efficiency and emission reduction have been critical research topics over several decades. Recent government regulations have shifted this focus to another important area related to engine health and performance monitoring. Although the advancements in the use of deep learning methods for system monitoring have shown promising results in this direction, designing efficient methods suitable for field systems remains an open research challenge. To this end, the objective of this study is to develop a computationally efficient neural network-based approach for identifying unknown parameters of a “mean value” diesel engine model to facilitate physics-based health monitoring and maintenance forecasting. We propose a hybrid method combining physics-informed neural networks (PINNs) and a deep neural operator (DeepONet) to predict unknown parameters and gas flow dynamics in a diesel engine. The operator network predicts independent actuator dynamics learnt through offline training, thereby reducing the PINN’s online computational cost. Additionally, empirical relationships in the mean-value model are represented using deep neural networks trained offline. To address PINN’s need for retraining with changing input scenarios (condition of the ODE/PDE or measured data), we propose two transfer learning (TL) strategies. The first strategy involves multi-stage transfer learning for parameter identification. While this method is computationally efficient as compared to online PINN training, improvements are required to meet field requirements. The second TL strategy focuses solely on training the output weights and biases of a sub-set of multi-head networks pre-trained on a larger dataset, substantially reducing computation time during online prediction. We also evaluate our model for epistemic and aleatoric uncertainty by incorporating dropout in pre-trained networks and Gaussian noise in the training dataset. This strategy offers a tailored, computationally inexpensive, and physics-based approach for parameter identification in diesel engine sub-systems for health monitoring and diagnosis.
Keywords DeepONet  $\cdot$ Inverse problem  $\cdot$ Transfer learning  $\cdot$ Few shot learning  $\cdot$ Digital twin  $\cdot$ System-of-Systems
1 Introduction
The health monitoring of internal combustion engines has evolved significantly over the past decades, driven by the need for improved performance, reliability, and maintenance efficiency. Historically, manufacturers developed various diagnostics and service tools to meet customer, operational, and service needs. However, traditional monitoring approaches often face challenges such as lack of fault isolation, limited prognostic capabilities, and robustness issues due to system variability and gradual degradation. For instance, system-level monitors may indicate a fault due to multiple slightly degraded components, none of which are individually degraded enough to trigger a fault on their own. Such scenarios frequently lead to trial-and-error methods for fault resolution and multiple service visits for end-users. Early strategies to address these challenges included the implementation of component-level monitoring with very tight thresholds. However, this approach proved problematic due to the increased risk of false positives and the challenges associated with developing and tuning high-fidelity monitors for each component. Consequently, various alternative solutions have been explored, including the use of complex logic algorithms, multiple degradation thresholds, and out-of-mission system tests. The proliferation of numerous small diagnostic components designed to cover all aspects of system monitoring has resulted in a highly complex and costly landscape for development, tuning, documentation, and risk management.
The present study aims to streamline the development of health monitoring systems, improve fault isolation, system health prognostics, and predictive maintenance capabilities. Our focus is on advancing system-level monitoring solutions that utilize all available data to provide high-fidelity monitoring of system performance while also delivering robust fault isolation and prognostic capabilities. While the addition of electronics, sensors, and control systems helps in improving system performance, timely health monitoring of the engine’s sub-systems is crucial to ensure optimal performance under all operating conditions. The integration of a wide array of sensors and actuators has enabled the assimilation of accurate data during engine operation. The availability of sensory data has catalyzed active research in the use of neural networks for engine modeling and fault diagnosis. The complex non-linear mapping learned by neural networks offers significant advantages for practitioners, given the inherent challenges in creating accurate physics-based models for combustion engines. By evaluating the important design parameters of the engine from field data recorded by sensors, it is possible to infer the health of the engine by comparing them with the actual design parameters. Numerous studies have demonstrated the application of neural networks in modelling different aspects of diesel engine [1, 2, 3, 4, 5, 6, 7, 8, 9]. However, majority of these works focus on data-driven models that rely on data measurements to predict engine states. In an effort to integrate known physics with measurement data, [10] proposed a hybrid model for predicting gas exchange processes in a diesel engine for individual sub-systems such as intake and exhaust manifold. This work focuses primarily on estimating the gas flow states for individual sub-systems using physics and data driven models but does not explore the idea of parameter estimation for the overall engine model. Thus, there exists a research opportunity to integrate physics-based models with data-driven approaches for creating efficient surrogates that are suitable for field application.
As a step towards parameter-based health monitoring, we propose an integrated physics and data-based approach to identify unknown parameters of engine sub-systems by using a framework based on physics-informed neural networks [11] (PINNs). In particular, PINNs leverage the known physical laws, typically represented by differential equations, to train neural network parameters for predicting solutions to these equations. Moreover, PINNs can also be applied to predict the unknown parameters of a differential equation system using available output state data. This application mode is termed as the inverse analysis for parameter identification. The nature of these parameters depends on the characteristics of the problem. While PINNs have demonstrated success across a variety of applications [12, 13, 14, 15, 16, 17, 18, 19], their potential for fast prediction in physical systems remains an open research avenue. Thus, our research premise is as follows: given a set of measurements and a physics-based model, estimate the critical parameters for sub-systems that can help in engine health monitoring. In our previous study [20], we demonstrated how PINNs can be used for estimating unknown parameters and gas flow dynamics in a diesel engine model described in [21]. We showed that the proposed method could predict simultaneously both unknown parameters and dynamics accurately. However, a limitation of this work was the high computational cost for parameter estimation, primarily stemming from the need to re-train the PINN model every time inference was required on a new data signal. This limited the applicability of this framework in field systems that require rapid prediction capability. In this study, we propose a hybrid method combining PINNs with neural operator learning for accelerating parameter estimation. While PINNs is employed to estimate the state variable and unknown parameters, the states which are independent are approximated using a neural operator.
Specifically, we consider the Deep Operator Network (DeepONet) as our choice of the neural operator [22] and build on the work of [23], who used DeepONet for predicting the gas flow dynamics described by [21]. This DeepONet model proposed by [23] is not designed for estimating the parameters of critical sub-systems but provides rapid estimates for gas flow dynamics. Hence, our objective is to combine the fast prediction capability of the operator network with the hybrid PINN framework for accelerating parameter estimation on diesel engines. Furthermore, in order to further reduce the computational cost, we present two transfer learning (TL) approaches: a) multi-stage transfer learning, and b) a few-shot transfer learning using a larger multi-head pre-trained neural network. Pre-trained neural networks were considered to approximate empirical formulas in a previous study [20], and the results showed good approximation. One of the major advantages of considering a pre-trained neural network instead of the least squares method is that we do not need to assume the type of non-linearity between the input and the output. The neural network learns the non-linearity itself if sufficient data is provided in the training process. Tosun et al. [3] also observed that a neural network approximation is better than the least squares method. However, one of the major concerns in both neural network and least squares method approximation is how to incorporate the systematic (e.g., manufacturing) variability in parts of the engines in the empirical formulae, i.e., the engine considered to obtain data for the training of the pre-trained network may vary marginally from the engines used in field applications. In this study, we consider dropout in the pre-trained networks to account for such systematic variability.
In summary, we propose a hybrid deep operator-infused Physics-Informed Neural Network (PINN) method to reduce the computation time required for estimating unknown parameters on a diesel engine. The effectiveness and efficiency of the proposed method are validated through several examples. To further reduce the computation cost, we present two transfer learning approaches, a multi-stage transfer learning and a few-shot transfer learning method. The applicability and accuracy of the proposed methods are shown by considering specific examples. For this purpose, we generate training data for the inverse problem, utilizing the Simulink file [24] provided in [21]. We consider the input signal for generating our simulated data using field measurements from engines on a test bed. Additionally, we incorporate appropriate Gaussian noise in the simulated data for the inverse problem to mimic real-world field data and the expected aleatoric uncertainty. Detailed problem statements and different cases considered are discussed in section 1.2 along with a brief discussion on the engine model considered.
1.1 Main contributions
We summarize below the key contribution of the present study aimed at achieving rapid gas flow dynamics prediction and parameter estimation for diesel engine health monitoring.
  • We propose a system-of-systems approach for modeling diesel engines through neural networks by substituting each sub-systems of a diesel engine model with data and physics-based neural networks. Each of these neural networks work in tandem to replicate a complete mean value model of the diesel engine.
  • We propose an Operator-infused PINN framework for rapid estimation of unknown system parameters and dynamics in diesel engines. The operator network approximate the independent state variable while PINN is used to model the gas filling dynamics and estimate the unknown parameters.
  • We propose two hybrid transfer learning strategies to meet the computational requirements of field systems. These frameworks utilize transfer learning combined with pre-trained operator networks to meet computational and robustness requirements for field systems.
  • We employ a multi-head training approach utilizing a large corpus of prior training data. This approach also includes a derivative-free layer to calculate physics loss during transfer learning for improved computational efficiency. The multi-head model bears close association to a Foundation model [25] since it can be re-used for different scenarios.
  • We develop an efficient parameter estimation method suited for deployment on field systems. The robustness of our methods was tested by replicating real-world noisy conditions in simulated data.
We organize the rest of the article as follows: Section 2, covers the methodology used to solve the problem stated in section 1.2. We explain the basic concept of our approach in section 2.1 and provide detailed discussion in subsequent subsections. In section 2.3, we present the DeepONet and how we approximate EGR and VGT states. We introduce the hybrid model involving DeepONet and PINN in section 2.4. In section 2.5, we discuss a multi-stage transfer learning strategy followed by a discussion on a computationally efficient few-shot transfer learning strategy in section 2.6. We present the results for Pre-trained networks and DeepONets in section 3. In section 4, we present the results for hybrid-PINN and multi-stage transfer learning. The results of few-shot transfer learning are discussed in section 5. A comparison between the two transfer learning methods is discussed in section 6. Finally, a summary of the present study along with future directions are presented in section 7.
1.2 Problem setup
A mean value model for the gas flow dynamics of a diesel engine was proposed by [21], and we adopt this model for our present study. The main components of the engine are the intake manifold, the exhaust manifold, the cylinder, the EGR valve system, the compressor, and the turbine. The gas flow dynamics through various components of the engine is described by the state variables of its component and the dependent variables. Specifically, the model we consider [21] has eight states
$\bm{x}=\{p_{im},\;\;p_{em},\;\;X_{Oim},\;\;X_{Oem},\;\;\omega_{t},\;\;\tilde{u}_{egr1},\;\;\tilde{u}_{egr2},\;\;\tilde{u}_{vgt}\}$ (1)
where $p_{im}$ and $p_{em}$ are the intake and exhaust manifold pressure, respectively, $X_{Oim}$ and $X_{Oem}$ are the oxygen mass fractions in the intake and exhaust manifold, respectively, $\omega_{t}$ is the turbo speed. $\tilde{u}_{egr1}$ and $\tilde{u}_{egr2}$ are the two states for the EGR actuator dynamics, and $\tilde{u}_{vgt}$ represents the VGT actuator dynamics. The mean value engine model is then expressed as
$\dot{\bm{x}}=f(\bm{x},\bm{u},n_{e}),$ (2)
where $\bm{u}=\{u_{\delta},u_{egr},u_{vgt}\}$ is the control vector and $n_{e}$ is the engine speed; $u_{\delta}$ is the mass of injected fuel, $u_{egr}$ is the EGR actuator position and $u_{vgt}$ is the VGT valve position. A schematic for the engine model is shown in Fig. 1.
In this study, the states describing the oxygen mass fraction of the intake and exhaust manifold, i.e., $X_{Oim}$ and $X_{Oem}$, are not considered, as the rest of the states do not depend on these two states. Appendix A provides the governing equations that describe the complete engine model. Interested readers may also refer to [21] or our previous work [20] for more details. The governing equations for the six state variables considered in this work are as follows:
$\dfrac{d}{dt}p_{im}=\dfrac{R_{a}T_{im}}{V_{im}}(W_{c}+W_{egr}-W_{ei}),$ (3)
$\dfrac{d}{dt}p_{em}=\dfrac{R_{e}T_{em}}{V_{em}}(W_{eo}-W_{t}-W_{egr}),$ (4)
$\dfrac{d}{dt}\omega_{t}=\dfrac{P_{t}\eta_{m}-P_{c}}{J_{t}\omega_{t}},$ (5)
$\dfrac{d\tilde{u}_{egr1}}{dt}$ $=\dfrac{1}{\tau_{egr1}}\left[u_{egr}(t-\tau_{degr})-\tilde{u}_{egr1}\right],$ (6)
$\dfrac{d\tilde{u}_{egr2}}{dt}$ $=\dfrac{1}{\tau_{egr2}}\left[u_{egr}(t-\tau_{degr})-\tilde{u}_{egr2}\right],$ (7)
$\dfrac{d\tilde{u}_{vgt}}{dt}$ $=\dfrac{1}{\tau_{vgt}}\left[u_{vgt}(t-\tau_{dvgt})-\tilde{u}_{vgt}\right].$ (8)
The calculation of $T_{em}$ in Eq. (4) (see Eq. (A.15)) involves two additional equation Eqs. (9) and (10). Here, $T_{1}$ is the temperature when the inlet valve closes after the intake stroke and mixing, and $x_{r}$ is the residual gas fraction:
$T_{1}=x_{r}T_{e}+(1-x_{r})T_{im},$ (9)
$x_{r}=\dfrac{\Pi_{e}^{1/\gamma_{a}}x_{p}^{-1/\gamma_{a}}}{r_{c}x_{v}}.$ (10)
Fig. 1: A schematic diagram of the mean value diesel engine with a variable-geometry turbocharger and exhaust gas recirculation [21]. The main components of the engine are the intake manifold, the exhaust manifold, the cylinder, the EGR valve system, the compressor, and the turbine. The control input vector is $\bm{u}=\{u_{\delta},u_{egr},u_{vgt}\}$, and the engine speed is $n_{e}$. The gas flow rates through the different subsystems are denoted by $W_{i}$’s (in blue) and represent the gas flow dynamics. The blue flow path represents clean airflow while the red path represents air-combustion mixture. (Source: Figure adapted from [21])
In a real system, measurements corresponding to output states $p_{im}$, $p_{em}$, $\omega_{t}$, $W_{egr}$ and $T_{em}$ are made along with the control inputs ($\bm{u}$) and engine speed $n_{e}$ through a network of integrated onboard sensors. The intake ($p_{im}$) and exhaust ($p_{em}$) manifold pressures are measured using pressure sensors. The turbocharger shaft speed ($\omega_{t}$) is measured using a rotational speed sensor, while the EGR valve flow rate ($W_{egr}$) is measured using a flow venturi tube. A few critical parameters of the diesel engine that are difficult to estimate on a running engine include the maximum effective area of the EGR valve ($A_{egrmax}$), compensation factor for non-ideal cycles ($\eta_{sc}$), the total heat transfer coefficient of the exhaust pipes ($h_{tot}$), and the maximum area of VGT ($A_{vgtmax}$). By estimating the variation in values of these parameters during engine operation, one can assess the health of the particular sub-system, thereby supporting FDI.
The objectives of the present study are twofold: (1) to predict the gas flow dynamics of all relevant engine states using a physics-informed deep learning framework, and (2) to identify critical unknown parameters in the system, given field measurements of $p_{im}$, $p_{em}$, $\omega_{t}$, $W_{egr}$, and $T_{em}$ and system models represented by equation (3)-(10). We refer to this as the inverse parameter identification problem in this work. As discussed in section 1, we aim to formulate a computationally efficient method suitable for use as an on-board monitoring system. The training data required for this study is generated using the Simulink file [24] accompanied with [21]. This Simulink file solves the differential equations discussed above with input vector $\bm{u}$ and engine speed $n_{e}$. Although we consider simulated data for training our neural network model, the input data (i.e. $\bm{u}$ and $n_{e}$) used to generate the simulation results are collected from actual engine test bed, thus representing real-world conditions. Additionally, we assess the effect of aleatoric uncertainty on our model’s performance using Gaussian noise. Thus, we consider two cases for our study:
  1. 1.
    Clean data: simulated noise-free data,
  2. 2.
    Noisy data: simulated data where appropriate Gaussian noise is added to $p_{im}$, $p_{em}$, $\omega_{t}$, $W_{egr}$, and $T_{em}$ to represent noisy sensor measurements expected in the field.
Since the input vector $\bm{u}$ and engine speed $n_{e}$ are collected from actual engine running conditions, they are assumed to be inherently noisy and no further noise is added to the input signals.
2 Methodology
2.1 Main concept
We formulate a hybrid model combining DeepONet and PINN for the inverse problem described in the previous section. The method may be subdivided into four parts.
  1. 1.
    Empirical relationships: First, we consider various empirical relationships between variables in the engine model. In our previous study [20], we demonstrated how these empirical formulae can be approximated using deep neural networks (DNN). The empirical formulae are engine-specific, and DNNs can provide a generalized representation when the exact relationships are unknown by training the DNN with data from the specific engine. In addition, to account for the systematic (e.g. manufacturing) variations of the engine, we introduce stochasticity into the empirical relationship by introducing dropout in the DNN.
  2. 2.
    Independent States Approximation: Second, we note here that the states describing the EGR and VGT dynamics are independent of the rest of the states, with no unknown parameters in these dynamics. We approximated these states using two DeepONets, one for the two EGR states ($\tilde{u}_{egr1}$ and $\tilde{u}_{egr2}$) and one for the VGT state ($\tilde{u}_{vgt}$) as discussed in section 2.3. These DeepONets are pre-trained using simulation data and used as surrogates to predict the actuator states.
  3. 3.
    Hybrid method: Third, we propose a hybrid method combining the pre-trained DNN, the pre-trained DeepONet and PINN model, as detailed in section 2.4.
  4. 4.
    Transfer Learning Strategies: Lastly, to reduce the computation cost (in terms of the number of epochs and trainable parameters), we consider two transfer learning strategies:
    1. (a)
      a hybrid PINN multi-stage transfer learning method, discussed in section 2.5, and
    2. (b)
      a hybrid PINN few-shot transfer learning method, discussed in section 2.6.
    We also use self-adaptive weights [26] for better and faster convergence in all examples of the proposed methods.
Furthermore, as noted above, we have pre-trained networks for empirical formulae (DNNs) and for EGR and VGT states (DeepONets). These networks are trained using simulated data (we label this as laboratory data for nomenclature simplicity), which corresponds to the design value of the unknown parameters. However, in the field, the values of the unknown parameters may change due to natural systematic changes over time. Thus, we test our proposed hybrid PINN method on a set of parameters different from the design parameters used for generating laboratory data for training our empirical networks and DeepONets. This helps us validate whether our hybrid PINN model can accurately identify parameters independently of the predictions from the pre-trained networks, thus allowing us to analyze any potential bias originating from this hybrid approach. The different values of the unknown parameters considered for generating laboratory and field data are shown in Table 1.
Table 1: True value of estimated parameters considered for generating laboratory and field data conditions. The laboratory data is considered for training the DNNs for empirical formulae, DeepONets for EGR and VGT actuator dynamics, and training the body for the few-shot transfer learning. The field data represents data that is used for training the hybrid PINN framework for parameter estimation. For confidentiality of data, we have shown the true values for field data as 1, and the laboratory data are scaled with respect to the field data.
$A_{egramx}$ $A_{vgtmax}$ $\eta_{sc}$ $h_{tot}$ Remarks
Laboratory data $1.333$ $1.1663$ $0.9179$ $1$ Considered for data generation to train the pre-trained networks and DeepONet
Field data $1$ $1$ $1$ $1$ Considered for data generation for field data (use in PINN inverse problem)
Mask and scaling consider in PINN
Mask function $\exp(.)$ $\exp(.)$ softplus(.) $\exp(.)$ This ensures the unknown parameters are always positive
Scaling ×$\times 10^{-4}$$\times 10^{-4}$$\times 1$$\times 10$ Scale to obtain the parameters in physical domain
2.2 Neural network surrogates for empirical formulae
The empirical formulae such as the polynomial functions for the volumetric efficiency ($\eta_{vol}$), effective area ratio function for EGR valve ($f_{egr}$), turbine mechanical efficiency ($\eta_{tm}$), effective area ratio function for VGT ($f_{vgt}$), choking function (for VGT) ($f_{\Pi_{t}}$), compressor efficiency ($\eta_{c}$), and volumetric flow coefficient (for the compressor) ($\Phi_{c}$) proposed in the mean value engine model in [21] are approximated using neural networks in our study. Dropouts are used in the DNN models to account for systematic variations. In Table 2, we present the six DNNs that approximate the empirical formulae along with their input and output variables.
Table 2: Neural network surrogates for empirical formulae $\mathcal{N}_{i}^{(P)}(\bm{x};\bm{\theta}_{i}^{P})$, $i=1,\dots,6$ denotes the surrogate for the $i^{th}$ DNN parameterized by $\theta_{i}^{P}$ with input $\bm{x}$. All these networks have one output each.
Surrogate $\mathcal{N}_{1}^{(P)}(\bm{x};\bm{\theta}_{1}^{P})$ $\mathcal{N}_{2}^{(P)}(\bm{x};\bm{\theta}_{2}^{P})$ $\mathcal{N}_{3}^{(P)}(\bm{x};\bm{\theta}_{3}^{P})$ $\mathcal{N}_{4}^{(P)}(\bm{x};\bm{\theta}_{4}^{P})$ $\mathcal{N}_{5}^{(P)}(\bm{x};\bm{\theta}_{5}^{P})$ $\mathcal{N}_{6}^{(P)}(\bm{x};\bm{\theta}_{6}^{P})$
Input ($\bm{x}$) $\{p_{im},n_{e}\}$ $\tilde{u}_{egr}$$\{\tilde{u}_{vgt},\Pi_{t}\}$$\{\omega_{t},T_{em},\Pi_{t}\}$$\{W_{c},\Pi_{c}\}$$\{T_{amb},\Pi_{c},\omega_{t}\}$
Output $\eta_{vol}$ $f_{egr}$ $F_{vgt,\Pi_{t}}$ $\eta_{tm}$ $\eta_{c}$ $\Phi_{c}$
Output restriction $S(f_{egr})$ $1.1\times S(F_{vgt,\Pi_{t}})$ $\text{min}(0.818,\eta_{tm})$ $\text{max}(0.2,S(\eta_{c}))$ $S(\Phi_{c})$
Equation A.6 A.23 A.31 A.33 A.40 A.50
$S\longrightarrow$ sigmoid function
The empirical equations are discussed in Appendix A
The DNNs for the empirical formulae are trained using a large laboratory dataset that captures all possible operating scenarios for the engine. We consider a mean square loss function for the training of these networks,
$\mathcal{L}_{i}(\bm{\theta}_{i}^{P})$ $=\dfrac{1}{n_{i}}\sum_{j=1}^{n_{i}}\left[y_{i}^{(j)}-\hat{y}_{i}^{(j)}\right]^{2},\;\;\;\;\;i=1,2,\dots,6$ (11a)
$=\dfrac{1}{n_{i}}\sum_{j=1}^{n_{i}}\left[y_{i}^{(j)}-\mathcal{N}_{i}^{(P)}(\bm{x}_{i};\bm{\theta}_{i}^{P})^{(j)}\right]^{2}$ (11b)
where the subscript $i=1,2,\dots,6$ indicate different neural networks shown in Table 2, $n_{i}$ represents the number of labelled training data for $i$th neural network. $\bm{x}_{i}$ denotes the input corresponding to the $i$th network, $\hat{y}_{i}$ and $y_{i}$ represent the estimated output and the corresponding labelled values, respectively for the $i$th network. The laboratory data used for training each neural network have been discussed in our previous work [20] along with detailed calculations for each labelled data. We also consider a 3% Gaussian noise in the simulated data of the output variables during training to account for the field measurement conditions. These networks are trained using the Adam optimizer. After training, we integrate these DNNs into our hybrid PINN framework to replace the empirical models. These pre-trained neural networks are shown with the output $g(\hat{y})$ in Fig. 2 and Fig. 4 in sections 2.4 and 2.6.
2.3 DeepONets for EGR and VGT states
Based on the universal approximation theorem of operators [27], Lu et al. [22] proposed the Deep Operator Network (DeepONet), which maps an infinite-dimensional function ($u\in\bm{U}$) to another infinite dimensional output function ($v\in\bm{V}$) in the Banach space. The conventional DeepONet consists of two DNNs, known as branch and trunk networks. The trunk network takes the spatio-temporal information (e.g. time $t$, domain points ${x,y,z}$) as input and provides the bases function as output. The branch network takes the input function ($u\in\bm{U}$) at fixed points and approximates the coefficients for each basis function. Importantly, the DeepONet architecture does not impose restrictions on the type of networks used for trunk and branch networks. The ability of DeepONet to learn complex mapping between input and output functional spaces makes it suitable for various applications in scientific machine learning. Mathematically, DeepONet operates as an operator that maps $\mathcal{G}_{\bm{\theta}}:\bm{U}\rightarrow\bm{V}$, and the output is given as
$v(t)\approx\mathcal{G}_{\theta}(u)(t)=\sum_{i=1}^{p}br(u)_{i}tr(t)_{i},$ (12)
where $br(u)_{i}$ and $tr(t)_{i}$ are the outputs of the branch and the trunk nets, respectively; $p$ is the number of neurons at the last layer of each of the branch and trunk networks, which corresponds to the number of base functions used in approximation of the spatio-temporal points. $u_{1},u_{2},\dots u_{n}$ are the input function sampled at $n$ fixed points. DeepONet has been applied to various engineering problems in its standard form or with appropriate modifications in its network and input structures. A comprehensive review and comparison of several variants can be found in [28]. As discussed in section 1, Kumar et al. [23] utilized DeepONet for predicting the gas flow dynamics of a diesel engine, albeit limited to accurate predictions of time sequences of up to five seconds. To extend the prediction length, the authors predict results every five seconds and stitch them to obtain long-duration results. In this case, the initial conditions are required every five seconds. Alternatively, the authors consider the predicted output of the DeepONet as the initial condition for the next time sequence. However, the authors in [23] observed that this approach results in higher error accumulation over time.
We note here that the engine is a continuous dynamical system, and this makes it challenging to identify proper operating sub-spaces for input-output mapping for the operator network learning. Hence, it is important to study the driving behaviour (engine running condition) over large duration and sample sizes. An appropriate window size (time duration) is required in which correct input-output function spaces can be assumed while preserving the dynamic behaviour of the system. Here, we consider a 60-sec window and approximate the independent state variables using Causality-DeepOnet [29], as proposed by Liu et al. [29]. Causality-DeepONet integrates causality and convolution of the input function in its formulation, which enhances accuracy in approximating oscillatory dynamic problems with a smaller training sample set. We find that this framework is particularly useful in capturing the rapidly changing dynamics of a diesel engine signal.
In this study, we use Causality-DeepONets to approximate the dynamics of the EGR actuator, i.e., $\tilde{u}_{egr1}$ (Eq. (6)) and $\tilde{u}_{egr2}$ (Eq. (7)) and the VGT actuator $\tilde{u}_{vgt}$ (Eq. (8)). Details of the two DeepONets are shown in Table 3. From hereon, we refer to Causality-DeepONet simply as DeepONet unless stated otherwise. These DeepONets approximate solutions to non-homogeneous initial value problems by learning the operator mapping between input function ($u$) and the initial condition ($\tilde{u}_{0}$) to the output ($\tilde{u}(t))$ using individual branch networks for the input function and initial condition, as described in Table 3. We note that the equations governing EGR and VGT actuator dynamics are first-order differential equations that can be computed efficiently using fast numerical methods. However, discrepancies between the mathematical model and experimental data were observed in [21]. By employing DeepONets trained on historical data, we obtain a better representation of real system dynamics as compared to a numerical method that relies on equations provided. Hence, our approach of using DeepONet surrogates for predicting actuator dynamics is likely to show better correlation when used on real system.
Table 3: DeepONet surrogates considered for approximation of EGR and VGT states. $\mathcal{D}_{i}(t,\bm{x};\bm{\theta}_{DP,i}),\,i=1,2$ denotes the DeepONet surrogate for the $i^{th}$ DeepONet parameterized by $\bm{\theta}_{DP,i}$. $\mathcal{D}_{1}(t,\bm{x};\bm{\theta}_{DP,1})$ approximates the two EGR states while $\mathcal{D}_{2}(t,\bm{x};\bm{\theta}_{DP,2})$ approximates the VGT state. The input to the trunk is time $t$ while $\bm{x}$ is the input to the branch. A total of four branches are used for $\mathcal{D}_{1}(t,\bm{x};\bm{\theta}_{DP,1})$ with two branches each taking $u_{degr}$ as input and the other two take initial condition $\{\tilde{u}_{egr1,0}$, $\tilde{u}_{egr2,0}\}$ as input.
DeepONet Variable Eq. No. of branch Branch inputs ($\bm{x}$)
$\mathcal{D}_{1}(t,\bm{x};\bm{\theta}_{DP1})$$\{\tilde{u}_{egr1}$, $\tilde{u}_{egr2}\}$ (A.16) and (A.17) 4 $\{u_{degr}$, $\tilde{u}_{egr1,0}\}$, $\{u_{degr}$, $\tilde{u}_{egr2,0}\}$
$\mathcal{D}_{2}(t,\bm{x};\bm{\theta}_{DP2})$$\{\tilde{u}_{vgt}\}$ (A.25) 2 $u_{dvgt}$, $\tilde{u}_{vgt,0}$,
Thus, we consider the operator mapped from the input function ($u$) and the initial condition ($\tilde{u}_{0}$) to the output ($\tilde{u}(t))$. We consider multiple branch nets for the input function and initial value as illustrated in Table 3. We also like to note that the equations describing the EGR and VGT states are first-order differential equations. Thus, it is possible to implement a fast numerical algorithm for their solutions. However, it was observed in [21] that the mathematical model and experimental data show slight disparity when an ensembled mean is considered. By considering DeepONet trained with historical data, we can obtain an average model of the output, which is a better representation than the numerical model.
$\mathcal{G}_{\bm{\theta}}:u,\tilde{u}_{0}\rightarrow\tilde{u}(u,\tilde{u}_{0})(t)$ (13)
We train these two DeepONets using labelled data generated using Simulink model, representing laboratory test measurements in real systems. For training of the EGR actuator states, labelled dynamic data of $\tilde{u}_{egr1}$ and $\tilde{u}_{egr2}$ are used, corresponding to inputs $u_{egr}$, $\tilde{u}_{egr1,0}$ and $\tilde{u}_{egr2,0}$. Similarly, for the training of VGT actuator states, labelled dynamic data $\tilde{u}_{vgt}$ for inputs $u_{vgt}$ and $\tilde{u}_{vgt,0}$ are used. The loss function for DeepONet training is formulated as:
$\mathcal{L}_{i}(\bm{\theta}_{DP,i})=$ $\dfrac{1}{n_{i}}\sum_{j=1}^{n_{i}}\left[y_{i}^{(j)}-\hat{y}_{i}^{(j)}\right]^{2},\;\;\;\;\;i=1,2$ (14a)
$=$ $\dfrac{1}{n_{i}}\sum_{j=1}^{n_{i}}\left[y_{i}^{(j)}-\mathcal{D}_{i}(t,\bm{x}_{i};\bm{\theta}_{DP,i})^{(j)}\right]^{2},\;\;\;\bm{x}\rightarrow\text{Branch inputs}$ (14b)
where $i=1,2$ denote the different DeepONets for approximating the EGR and VGT states; $\bm{x}_{i}$ represents the corresponding input to the $i^{th}$ DeepONet, $\hat{y}_{i}$ and $y_{i}$ are the output of the $i^{th}$ DeepONet, and the corresponding labelled data, respectively, and $n_{i}$ is the number of labelled dataset for the $i^{th}$ DeepONet.
We also note that the DeepONet for the EGR states has two output quantities. Consequently, the loss function for this DeepONet is the weighted summation of the loss function for individual output quantities. The network architecture, loss function and input-output are discussed in more detail in section 3, along with specific examples. These networks are trained using Adam optimizer. As discussed earlier, we utilize simulation data to train the DeepONets and subsequently replace the EGR and VGT dynamics with these pre-trained networks within the PINN model. Note that for DeepONet training here, we do not incorporate dropouts since the outputs of these DeepONets are used as inputs to the pre-trained empirical networks that include dropouts to inherently addresses the variability, as discussed in section 2.2. In the following section, we present details about our hybrid PINN framework to tackle the parameter estimation problem.
2.4 A hybrid model using DeepONet and PINN for the engine problem
In the previous section (2.3), we discussed how the three states corresponding to EGR and VGT can be approximated using DeepONet surrogates. In section 2.2, we introduced how we use DNNs to estimate empirical relationships in the engine model. In this section, we will discuss in detail how we construct a hybrid model combining DeepONet, empirical DNNs, and PINN to solve inverse problems. We propose a hybrid PINN model combining DeepONet and DNN to predict the gas flow dynamics and estimate unknown parameters.
Fig. 2 illustrates a schematic of the proposed hybrid model comprised of PINN and DeepONet. The first part of hybrid PINN model comprises a fully connected neural network and approximates the solution to a specific differential equation. The second part is the physics-informed segment, in which we encode the corresponding differential equation using automatic differentiation [30]. We denote the unknowns in the equation as $\bm{\Lambda}$. If the unknowns are constant, these are treated as trainable scalar variables. If the unknowns are fields, these are approximated using DNN(s). The empirical formulas are approximated using DNN as discussed in section 2.2. Additionally, the independent states related to actuator signals are approximated using DeepONets. These states (or the function of these states) may be incorporated into the loss function for PINN training. The loss function ($\mathcal{L}(\bm{\theta};\bm{\mathcal{\varLambda}})$ for solving the inverse problems consists of two components: the physics loss ($\mathcal{L}_{\text{phy}}(\bm{\theta};\bm{\mathcal{\varLambda}})$, and the data loss ($\mathcal{L}_{\text{data}}(\bm{\theta};\bm{\mathcal{\varLambda}})$)
$\mathcal{L}(\bm{\theta};\bm{\mathcal{\varLambda}})=\lambda_{1}\mathcal{L}_{\text{phy}}(\bm{\theta};\bm{\mathcal{\varLambda}})+\lambda_{2}\mathcal{L}_{\text{data}}(\bm{\theta};\bm{\mathcal{\varLambda}})$ (15)
where $\bm{\theta}$ denotes the set of parameters of the neural network, i.e. the weights, $\bm{W}$ and the biases, $\bm{b}$; $\bm{\mathcal{\varLambda}}$ are the unknown scalar parameters to be estimated. The parameters $\lambda_{1}$ and $\lambda_{2}$ are the weights which control the convergence of each loss component. These $\lambda$s values may be constant or self-adaptive [26] depending on the problem setup and solution method. The data loss may be considered as the mean squared error (MSE) loss of the known labelled data, given by:
$\mathcal{L}_{\text{data}}(\bm{\theta};\bm{\mathcal{\varLambda}})=\frac{1}{M}\sum^{M}_{i=1}\left[\hat{y}(t_{i};\bm{\theta})-y(t_{i})\right]^{2},$ (16)
where $M$ represents the number of measured points, $\hat{y}(t_{i};\bm{\theta})$ and $y(t_{i})$ denote the prediction of neural network and ground truth data, respectively, at time $t_{i}$. For multiple quantities, the data loss may be considered as the summation of individual losses for each quantity. The physics loss may be written as
$\mathcal{L}_{\text{phy}}(\bm{\theta};\bm{\mathcal{\varLambda}})=\frac{1}{N}\sum^{N}_{i=1}\left[r(t_{i};\bm{\theta};\bm{\mathcal{\varLambda}})\right]^{2},$ (17)
where $N$ is the number of collocation points chosen for solving the equation and $r(t_{i};\bm{\theta};\bm{\mathcal{\varLambda}})$ represents the residual of the corresponding differential equation. The residual $r(t_{i};\bm{\theta};\bm{\mathcal{\varLambda}})$ should ideally be zero throughout the entire domain. The objective, therefore, is to obtain the optimal parameters ($\bm{\theta}$) of the neural network as well as the unknown parameters ($\bm{\mathcal{\varLambda}}$) in the system. We minimize the loss function Eq. (15) to obtain the optimal values of these parameters. Furthermore, we also incorporate self-adaptive weights [26] in the loss. We minimize the loss function with respect to the parameters ($\bm{\theta}$ and $\bm{\varLambda}$) and maximize with respect to the self-adaptive weights ($\lambda$). Thus, the optimization process may be written as,
$\min\limits_{\bm{\theta},\bm{\varLambda}}\;\max\limits_{\lambda}\mathcal{L}(\bm{\theta},\varLambda)$ (18)
If we consider a gradient descent-ascent approach, we can express the optimization process as
$\bm{\theta},\bm{\varLambda}$ $=\bm{\theta},\bm{\varLambda}-lr_{\theta,\varLambda}\nabla_{\theta,\varLambda}\mathcal{L}(\bm{\theta},\varLambda)$ (19a)
$\lambda$ $=\lambda+lr_{\lambda}\nabla_{\lambda}\mathcal{L}(\bm{\theta},\varLambda),$ (19b)
where $lr_{\theta}$ denotes the network optimiser learning rate for $\bm{\theta}$, $lr_{\varLambda}$ represents the learning rate for the optimizer used to identify unknown scalar parameters $\bm{\varLambda}$, and $lr_{\lambda}$ refers to the learning rate for the optimizer used for updating self-adaptive weights. Incorporating distinct learning rates and optimizers allows for fine tuning of each component in this hybrid framework for efficient convergence.
To integrate the complete hybrid model, we combine the DNN models used for approximating the empirical formulae as discussed in section 2.2 and the two DeepONet models for approximating actuator states $\tilde{u}_{egr1}$, $\tilde{u}_{egr2}$ and $\tilde{u}_{vgt}$ with the PINN DNN model, as shown in figure 2. The parameters of these DeepONets and DNNs (for the empirical formulae) are kept fixed when we train the hybrid PINN model. Furthermore, in Eq. (15), we demonstrated how PINNs can be used for solving a single differential equation. For systems involving multiple equations, one can either use one DNN with multiple outputs or multiple DNNs each with one output. The loss function in Eq. (15) can be extended to include the loss function for all the equations and data by summing up individual loss functions for each equation and data. This hybrid approach leverages pre-trained networks and a PINN network that can be trained on live data for efficient and accurate estimation of engine parameters and state variables.
Fig. 2: Schematic of a hybrid model of DeepONet and physics-informed neural networks (PINNs) for inverse problems. In the left part of the figure (enclosed in red dashed line), a PINN approximates the solution $y$ to a differential equation with time $t$ as input. The top left section (enclosed within black hatched lines) represents another DNN which takes as input $\hat{y}$ (and potentially other inputs such as ambient condition) and outputs a function $g(\hat{y})$. This network is pre-trained with labelled laboratory data. On the top right section of the figure (enclosed in black dashed line), a DeepONet model takes inputs of time $t$ and an input function $u$ to approximate a state $G(u)(t)$. This network is also pre-trained with labelled laboratory data. The physics/residual loss is shown in the bottom right part of the figure (enclosed within the blue dashed line). The PINN approximates the solution to differential equations. The total loss $\mathcal{L}(\bm{\theta})$ includes the loss of equation as well as the data. Weights $\lambda_{1}$ and $\lambda_{2}$ are applied to the data and physics loss respectively, and may be fixed or adaptive depending on the problem and solution method. $\bm{\theta}=\{\bm{W},\bm{b},\bm{\Lambda}\}$ represents the parameters of the PINN, where $\bm{W}$ and $\bm{b}$ are the weights and biases, and $\bm{\Lambda}$ are the unknown parameters of the ODE; $\sigma$ denotes the activation function used for the PINN model, and $r$ is the residual for the equation. Further, $\bm{\theta}^{P}=\{\bm{W}^{P},\bm{b}^{P}\}$ denotes the parameters of the pre-trained neural network $g(y)$. Similarly, $\bm{\theta}_{DP}=\{\bm{W}_{DP},\bm{b}_{DP}\}$ represents the parameters of the pre-trained DeepONet $G(u)(t)$.
In section 1.2, we discussed the inverse problem of interest in this study. Our objectives are to: (1) learn the dynamics of the six gas flow dynamics state estimates, three of which are approximated using two DeepONets, and (2) infer the unknown parameters in the system, given field measurements of system variables $\{p_{im},p_{em},\omega_{t},W_{egr},T_{em}\}$ as well as Eqs. (3)-(10), using PINNs. To achieve these objective, we utilize four DNNs as surrogates for the solutions to different equations, as shown in Table 4. We encode Eqs (3)-(10) using the automatic differentiation and formulate the physics-informed loss. The total loss function considered for training the PINNs model is given by:
$\begin{split}\mathcal{L}(\bm{\theta},\bm{\mathcal{\varLambda}},\bm{\lambda}_{p_{im}},\bm{\lambda}_{p_{em}},\bm{\lambda}_{\omega_{t}},\bm{\lambda}_{W_{egr}},\lambda_{T_{em}},\lambda_{T_{1}})=&\mathcal{L}_{p_{im}}+\mathcal{L}_{p_{em}}+\mathcal{L}_{\omega_{t}}+100\times\mathcal{L}_{x_{r}}+\lambda_{T_{1}}\times\mathcal{L}_{T_{1}}+\\ &\mathcal{L}^{ini}_{p_{im}}+\mathcal{L}^{ini}_{p_{em}}+\mathcal{L}^{ini}_{\omega_{t}}+100\times\mathcal{L}^{ini}_{x_{r}}+100\times\mathcal{L}^{ini}_{T_{1}}+\\ &\mathcal{L}^{data}_{p_{im}}(\bm{\lambda}_{p_{im}})+\mathcal{L}^{data}_{p_{em}}(\bm{\lambda}_{p_{em}})+\mathcal{L}^{data}_{\omega_{t}}(\bm{\lambda}_{\omega_{t}})+\\ &\mathcal{L}^{data}_{W_{egr}}(\bm{\lambda}_{W_{egr}})+\lambda_{Tm}\mathcal{L}^{data}_{T_{em}},\end{split}$ (20)
Table 4: Neural network for PINN: Neural network surrogates considered to approximate the state variables and $x_{r}$ and $T_{1}$ for solving the inverse problems. $\mathcal{N}_{i}(t;\bm{\theta}_{i}),i=1,...,4$ denotes the surrogate for the $i^{th}$ DNN parameterized by $\bm{\theta}_{i}=\{\bm{W},\bm{b}\}_{i}$. The input to these DNN is time $t$. DNN $\mathcal{N}_{1}(t;\bm{\theta}_{1})$ has two outputs for $\{p_{im}$ and $p_{em}\}$; the remaining DNNs have only one output.
Surrogate $\mathcal{N}_{1}(t;\bm{\theta}_{1})$ $\mathcal{N}_{2}(t;\bm{\theta}_{2})$ $\mathcal{N}_{3}(t;\bm{\theta}_{3})$ $\mathcal{N}_{4}(t;\bm{\theta}_{5})$
Output Variables ${p}_{im}$, ${p}_{em}$ ${x}_{r}$ ${T}_{1}$ ${\omega}_{t}$
Output transformation $S_{p}(p_{em})$, $S_{p}(p_{im})$ $S(x_{r})$ $S_{p}(T_{1})+230./300.0$ $S_{p}(\omega_{t}$)
Output scaling ×$\times 10^{5}$$\times 1$$\times 300$$\times 5.0\times 10^{3}$
Equation (3) and (4) (10) (9) (5)
   $S_{p}\rightarrow$ Softplus(),    $S\rightarrow$ sigmoid()
The individual physics losses are represented as:
$\mathcal{L}_{p_{im}}=$ $\dfrac{1}{n}\sum_{k=1}^{n}r(p_{im}^{(k)})^{2}=\dfrac{1}{n}\sum_{k=1}^{n}\left(\dfrac{dp_{im}^{(k)}}{dt}-\dfrac{R_{a}T_{im}}{V_{im}}\left(W_{c}^{(k)}+W_{egr}^{(k)}-W_{ei}^{(k)}\right)\right)^{2}$ (21a)
$\mathcal{L}_{p_{em}}=$ $\dfrac{1}{n}\sum_{k=1}^{n}r(p_{em}^{(k)})^{2}=\dfrac{1}{n}\sum_{k=1}^{n}\left(\dfrac{dp_{em}^{(k)}}{dt}-\dfrac{R_{e}T_{em}^{(k)}}{V_{em}}\left(W_{eo}^{(k)}-W_{t}^{(k)}-W_{egr}^{(k)}\right)\right)^{2}$ (21b)
$\mathcal{L}_{\omega_{t}}=$ $\dfrac{1}{n}\sum_{k=1}^{n}r(\omega_{t}^{(k)})^{2}=\dfrac{1}{n}\sum_{k=1}^{n}\left(\dfrac{d\omega_{t}^{(k)}}{dt}-\dfrac{P_{t}^{(k)}\eta_{m}-P_{c}^{(k)}}{J_{t}\omega_{t}^{(k)}}\right)^{2}$ (21c)
$\mathcal{L}_{x_{r}}=$ $\dfrac{1}{n}\sum_{k=1}^{n}r(x_{r}^{(k)})^{2}=\dfrac{1}{n}\sum_{k=1}^{n}\left(x_{r}^{(k)}-\dfrac{(\Pi_{e}^{(k)})^{1/\gamma_{a}}(x_{p}^{(k)})^{-1/\gamma_{a}}}{r_{c}x_{v}^{(k)}}\right)^{2}$ (21d)
$\mathcal{L}_{T_{1}}=$ $\dfrac{1}{n}\sum_{k=1}^{n}r(T_{1}^{(k)})^{2}=\dfrac{1}{n}\sum_{k=1}^{n}\left(T_{1}^{(k)}-\left(x_{r}^{(k)}T_{e}^{(k)}+(1-x_{r}^{(k)})T_{im}\right)\right)^{2}$ (21e)
where $n$ and $r(.)$ are the number of residual points and residual, respectively. The data losses for the known field data are given by:
$\mathcal{L}^{data}_{\phi}(\bm{\lambda}_{\phi})=$ $\dfrac{1}{n_{\phi}}\sum_{j=1}^{n_{\phi}}\left[\left(\phi_{data}^{(j)}-\hat{\phi}_{data}^{(j)}\right)\lambda_{\phi}^{(j)}\right]^{2},\quad\quad\phi=p_{im},p_{em},\omega_{t},W_{egr}$ (22a)
$\mathcal{L}^{data}_{T_{em}}=$ $\dfrac{1}{n_{T_{em}}}\sum_{j=1}^{n_{T_{em}}}\left[\left(T_{em_{data}}^{(j)}-\hat{T}_{em_{data}}^{(j)}\right)\right]^{2}$ (22b)
where $\{\phi,T_{em}\}$ and $\{\hat{\phi},\hat{T}_{em}\}$ are the true and predicted values of the variable, respectively, and $n_{\phi}$ is the number of data points for the variable. $\lambda$ is the self-adaptive weight associated with the variables.
The initial condition losses are given as:
$\mathcal{L}^{ini}_{\phi}=\dfrac{1}{1}\sum_{j=1}^{1}\left(\phi_{0}^{(j)}-\hat{\phi}_{0}^{(j)}\right)^{2},\;\;\;\;\;\;\;\;\phi=p_{im},p_{em},\omega_{t},x_{r},T_{1}$ (23)
The parameters of the neural networks for the hybrid PINN model (ref. Table 4), along with unknown parameters (ref. Table 1) and the self-adaptive weights ($\lambda$), are obtained by optimizing the loss function (20) using the Adam [31] optimizer. The optimization process occurs in two phases. In the first phase,we optimize the PINN network parameters, unknown parameters, and self-adaptive weights over a certain number of epochs. At the end of this phase, we freeze the values of the self-adaptive weights. Thereafter, we optimize only the parameters of the hybrid PINN model and the unknown parameters using Adam, keeping the self-adaptive weight constant with the values obtained at the end of the first phase. As discussed earlier, the parameters of the pre-trained neural networks for the empirical formula and the DeepONets are kept frozen during the hybrid PINN optimization process (both phases).
2.5 Multi-stage Transfer learning
In the preceding sections, we have outlined a method for predicting unknown parameters and gas flow dynamics of a diesel engine. In this method, the parameters of the networks, unknown parameters and the self-adaptive weights are initialized randomly and trained to obtain the optimal parameters. Due to the stochastic initialization of the network parameters, unknown parameters and self-adaptive weights, significant computation time is required to converge to an accurate estimate. Furthermore, any changes in the control input vector $u=\{u_{egr},u_{vgt},u_{\delta}\}$, engine speed $n_{e}$ or ambient conditions necessitates re-training the entire PINN model.
To mitigate the challenge related to computational time, we implement a multi-stage transfer learning to train the hybrid PINN model separately for each one-minute time segment as shown in figure 3. Transfer learning [32, 33, 34] utilizes knowledge gained from training a model on one dataset (typically referred to as source data) to enable knowledge transfer to train on a different but similar dataset (known as target data). We demonstrate our methodology using a five-minute long data signal divided into five one-minute segments. First, the complete hybrid PINN model is trained with the given control vector $\{u_{\delta},u_{egr},u_{vgt}$}, and engine speed $n_{e}$ for time segment $0-60$s. The weights and biases of the network, denoted by $\bm{\theta}$, are initialized randomly. The unknown parameters for the engine, $\bm{\Lambda}:=\{A_{egrmax},A_{vgtmax},\eta_{sc},h_{tot}\},\bm{\Lambda}\in\mathbb{R}^{+}$ are initialized with a random positive value (discussed in detail in section 4). Similarly, the self-adaptive weights $\bm{\lambda}$ are also initialized randomly at the beginning of the training. Thereafter, the PINN network is trained as per the process outlined in section 2.4, and the network parameters and unknown parameters $\bm{\Lambda}$ are saved. In the transfer learning stage, these saved parameter values are used for initialization of network parameter (weights and biases) and unknown parameters, i.e. instead of random initialization, we import these values from the trained model in $0-60$ sec. In Figure 3, we illustrate different stages of the proposed transfer learning scheme, and these stages include:
  1. Stage I:
    Transfer weights and biases ($\bm{\theta}$) and parameter estimated ($\bm{\Lambda}$) from the trained model on segment 0-60 sec to the new training in new time-segment. Initialize self-adaptive weights ($\bm{\lambda}$) randomly. Train complete PINN model and update $\bm{\theta},\bm{\Lambda}$, $\bm{\lambda}$} upto $N_{1}$ epochs.
  2. Stage II:
    Freeze all network weights ($\bm{\theta_{l-1}}$) except the last layer ($\bm{\theta_{l}}$). Update $\bm{\theta_{l}},\bm{\Lambda}$, $\bm{\lambda}$} upto $N_{2}$ epochs.
  3. Stage III:
    Freeze $\bm{\theta_{l-1}}$ and $\bm{\lambda}$, update $\bm{\theta_{l}},\bm{\Lambda}$} upto $N_{3}$ epochs.
Fig. 3: Schematic for Multi-stage transfer learning showing the network training protocol for different segments of time. In this approach, the hybrid PINN model is trained for the first segment of time (0-60 sec). Thereafter, for the subsequent time segments (60-300s in each one-minute segment), we transfer the network parameters ($\bm{\theta}$) and engine parameter estimated ($\bm{\Lambda}$) from the hybrid PINN model trained on the first time segment. Each segment uses a different control vector $\{u_{\delta},u_{egr},u_{vgt}$}, and engine speed $n_{e}$. The adaptive weights ($\bm{\lambda}$) are initialized at the start of training at epoch=0. The network weights and biases of all the layers ($\bm{\theta}$), self-adaptive weights ($\bm{\lambda}$), and unknown parameters ($\bm{\Lambda}$) are updated in Stage I training phase. In Stage II, only weights and biases of the last layer of the PINN model ($\bm{\theta_{l}}$) are trained while freezing all the hidden layers ($\bm{\theta_{l-1}}$). Additionally, $\bm{\lambda}$ and $\bm{\Lambda}$ continue to be updated in this phase. In Stage III, we continue to update $\bm{\theta_{l}}$ and $\bm{\Lambda}$ while the self-adaptive weights $\bm{\lambda}$ are fixed to the values obtained at the end of Stage II. This approach avoids the need to re-train the complete model again and helps reduce the computation time for parameter estimation. In total, for predicting across a 5-minute time window, we train five models, with the model trained for segment 1 (0-60 sec) serving as the reference model from which the remaining models acquire knowledge during the transfer learning phase.
2.6 Few-shot transfer learning
In the previous sections, first we developed a hybrid DeepONet and PINN model and proposed a multistage transfer learning for the engine inverse problem discussed in section 1.2. In the method, for the first time segment, the model needs to be trained from the beginning with random initialization of the weights and biases. Secondly, to calculate the physics loss, the derivative for the output of the neural network with respect to the input needs to be calculated at each epoch. Thus, although transfer learning improves the computation time, it is still high as per application requirements. One approach to improve computational time is by reducing (or eliminating) the differentiation operation involved in PINN training . This may be achieved by considering the derivative to be a linear combination of some base vector. For example, in function connection [35], multi-head network [36] or reservoir computing [37], we generally do not train the hidden layer and approximate the derivative of the output as a linear combination of the derivative of the hidden layer. Desai et al. [36] incorporated one-shot inference on new tasks by training the PINN network on a training task set and then freezing all except the last layer that is trained for the new task. This approach enables one to obtain the network output as a linear matrix multiplication of the output from the frozen layer with the weight matrix of the last layer. This reduces the need for automatic differentiation to evaluate the gradients of outputs with respect to inputs, which is computationally expensive. We adopt this transfer learning paradigm here, called few-shot transfer learning.
Fig. 4 illustrates a schematic diagram of a few-shot transfer learning combined with pre-trained DNN and DeepONet. Here, we note that the objectives of the pre-trained DNN and DeepONet are the same as discussed in the previous section. Hybrid PINN, using a few shot transfer learning, may be divided into two phases. The first phase (Phase I) is an offline training phase with multiple heads for unknown variables. In this phase, we trained the network with multiple heads, each head approximating one minute (corresponding to a different time zone) of the same output variable. The input to the network is time $0-60$ sec, normalized between $[-1,1]$. The first phase of offline training aims to obtain good universal basis functions (hidden layers), which can be combined to obtain appropriate output in Phase II for solving the inverse problem using PINN. In the first phase, the network is trained in a data-driven setting in an offline mode. Once the network is trained, the hidden layers (universal basis function) are transferred to the hybrid PINN model to estimate the variables in Phase II.
In Phase II, we transfer the hidden layer of the multi-head network trained in the first phase. We randomly initialize the last layer ($\bm{W}_{out},b_{out}$) and setup the hybrid PINN inverse problem for one minute duration independently. In this phase, only the last layer of the network and the unknown parameters of the equation and the self-adaptive weights are optimized. The estimated output is obtained as
$y(t)=\bm{h}(t)\bm{W}_{out}+b_{out},$ (24)
where $\bm{h}(t)$ is the output of the last hidden layer, $\bm{W}_{out}$ and $b_{out}$ are the weights and biases of the output layer. The derivative of the output w.r.t. the input variable is
$\dfrac{dy(t)}{dt}=\dfrac{d\bm{h}(t)}{dt}\bm{W}_{out}.$ (25)
Since in Phase II, the hidden layers are fixed (not trained), the derivative $\dfrac{d\bm{h}(t)}{dt}$ does not change and can be kept outside the training loop. Thus, the derivatives need to be calculated only once outside the training loop. In other words, we approximate the derivative of the output as a linear combination of the derivative of the hidden layer with the same combination matrix ($\bm{W}_{out}$) as that of the output. The loss function for the few shot transfer learning is the same as that of hybrid PINN given by Eq. (20). Note here that we utilize an output transformation function and hence we have to employ the chain rule for calculating the derivative of output with respect to input. We discuss this in detail in the numerical result section 5.
Fig. 4: Schematic diagram for the few-shot transfer learning architecture with multi-head network. The left side of the plot shows Phase I of the few-shot transfer learning approach. In this phase, we train the multi-head networks using laboratory data where each head approximates one minute of response (corresponding to a different time zone). Input to this network is time $0-60$ sec, normalized between $[-1,1]$. Phase I training is a data-driven offline training. On the right side of the plot, Phase II that involves transfer learning is shown. In this phase, we transfer the hidden layer and derivative of the output of the last hidden layer from the network trained in Phase I. We randomly initialize the weights and biases of the last layer along with the unknown parameters. The output and the derivative of the output are predicted using Eqs. 24 and 25, respectively. We formulate the hybrid PINN model by combining the PINN network with DeepONet and empirical DNN. The total loss ($\mathcal{L}_{\bm{\theta}}$) is calculated as the weighted sum of the physics and data loss. The weights $\lambda_{1}$ and $\lambda_{2}$ may be constant or adaptive depending on the problem setup and solution method. The trainable parameters are the weights and bias of the output layer of the PINN network ($\bm{\theta}_{out}=\{\bm{W},b\}$), the unknown parameters of the equation, the self-adaptive (in case of self-adaptive) weights. The weight and biases of the hidden layer of the PINN network, the parameters of the pre-trained network for empirical formula, and the parameters of the DeepONets remain fixed during the optimization process.
2.7 Data generation
As discussed earlier, we utilize two types of networks: (1) pre-trained DNN for the empirical formula and DeepONets (pre-trained) for the independent states ($\tilde{u}_{egr1}$, $\tilde{u}_{egr2}$ and $\tilde{u}_{vgt}$), (2) PINN networks used for online learning. Thus, two distinct types of data is required. We label the data used for training the empirical formulae networks and DeepONets for EGR and VGT states as laboratory data. The data used for solving the inverse problem through PINNs is labeled as field data, which is used to simulate a real testing scenario. Our prior study [20] provides an in-depth discussion on the laboratory data required for training each DNN for empirical formulae and other necessary details. In the absence of laboratory and/or field data, we use Simulink [24] to generate both datasets under different conditions as presented in Table 1.
In the present study, we consider four sets of input data (control vector and speed) to generate lab and field data as shown in Table 5. The simulated data are generated with $dt=0.2$ sec for training and testing of DNN for empirical formulae and DeepONet except training of $\eta_{tm}$. In the case of $\eta_{tm}$, these are generated at $dt=0.025$ sec. We trained the DNN for $f_{egr}(\mathcal{N}_{2}^{P}(:,\bm{\theta}))$ only with data from Case-I shown in table 5. The data for training of multi-head DNN for few-shot transfer learning are discussed later in the numerical results section. For further details regarding data generation, we refer readers to our previous study [20]. We consider PINN for each 60 sec. time segment independently in our computational study. The initial conditions for all the state variables and $x_{r}$ and $T_{1}$ are known for each time segment considered in the numerical study.
Table 5: Data considered for training and testing of pre-trained networks, DeepONets and field data for PINN. The input data Set-I is two hours of input control vector $\{u_{\delta},u_{egr},u_{vgt}\}$ and $n_{e}$. Set-II is twenty-minute of input control vector $\{u_{\delta},u_{egr},u_{vgt}\}$ and $n_{e}$. Similarly, Set-III twenty-four minute data marginally modified $u_{egr}$ to get a larger operating range. Similarly, Set-VI is a twenty-four-minute data marginally modified $u_{vgt}$ to get a larger operating range. Case-I to Case-IV are considered for the training of the surrogate neural network for the empirical formulae ($\mathcal{N}_{i}^{(P)}(:,\bm{\theta}_{i})$). Case-I and Case-VI are considered for the training, while Case-V is considered for the testing of these networks. The data for the field data are also considered from Case-V. Case-1 and Case-VI are considered for the training of DeepONet for EGR. Case-1 and Case-VII are considered for the training of DeepONet for VGT. DeepONet for EGR and VGT are tested using Case-V. $T$ and $P$ refer to the temperature and pressure of baseline measurement.
Case $T_{amb}$ $p_{amb}\times 10^{5}$ Input Sampling Purpose
(K) (Pa) $dt$
Case-I $T-65$ $P-0.100$ Set-I 1 sec Training pre-trained DNN for empirical formulae
Case-II $T-65$ $P+0.211$ Set-I 1 sec
Case-III $T-28$ $P-0.100$ Set-I 1 sec
Case-IV $T+15$ $P+0.211$ Set-I 1 sec
Case-V $T$ $P$ Set-II 0.2 sec
Testing of DNN and DeepONet / Field data for PINN
Case-I $T-65$ $P-0.100$ Set-I 1 sec Training $\tilde{u}_{egr1}$ and $\tilde{u}_{egr2}$
Case-VI Set-III 1 sec
Case-I $T-65$ $P-0.100$ Set-I 1 sec Training $\tilde{u}_{vgt}$
Case-VII Set-VI 1 sec
3 Results: Pre-trained networks and DeepONets
To enhance the organization and readability of our numerical results, we divide the discussion into three sections. First, we discuss the results of pre-trained DNNs used for estimating empirical relationships and the DeepONets for the EGR and VGT states. These two networks are integrated into the hybrid PINN framework. In section 4, we present the results pertaining to the application of PINNs and the multi-step transfer learning approach. In Section 5, we discuss numerical results for the few-shot transfer learning approach.
3.1 Pre-trained networks: Neural network for empirical formulae
The empirical formulae of the engine model are approximated using surrogate neural networks with dropout as discussed in section 2.2. We apply a 2% dropout for the hidden layers of these networks. Notably, dropout is employed not only during the training of these networks but also during the PINN optimization process for the inverse problem and prediction. This helps us to analyze the potential variation in PINN estimates expected due to systematic (e.g. manufacturing) variability for real systems. We also added a 3% Gaussian noise to the output state variables during training to replicate the expected noise in laboratory measurement.
3.2 DeepONet: Approximation of EGR states ($\tilde{u}_{egr1}$ and $\tilde{u}_{egr2}$) and VGT states ($\tilde{u}_{vgt}$) using DeepONets
[…]
The loss function for individual output variables is the mean square error between the true value ($\tilde{u}_{1}$) and the predicted one ($\tilde{u}_{1}(\bm{\theta})$). The total loss is the weighted sum of losses for both outputs. For EGR actuator dynamics, we use weights $a_{1}=1$ and $a_{2}=100$ (ref. Fig. 5). The DeepONets are trained using the Adam optimizer, with $3\times 10^{5}$ epochs for DeepONet $\mathcal{D}_{1}$ and $4\times 10^{5}$ epochs for DeepONet $\mathcal{D}_{2}$. The input and output variables are scaled and are as shown in Table 6 along with including the relative $L_{2}$ error for prediction. The mean and standard deviation of the error are calculated from six independent training sessions and show good accuracy. To further validate the model’s robustness, we added a $3\%$ Gaussian noise to the labelled output data and trained the networks with this noisy data. The relative $L_{2}$ error for these cases is also shown in Table 6, and we note only a marginal change in error, demonstrating the model’s robustness to noise in training data.
Fig. 5: Schematic diagram of DeepONet for prediction of EGR states showing branches and trunk. The trunk network takes time as its input, at which the value of the output function needs to be evaluated. The four branches take $u_{d}$, $\tilde{u}_{1,0}$, $u_{d}$ and $\tilde{u}_{2,0}$ as input and approximate two outputs $\tilde{u}_{1}=\tilde{u}_{egr1}$ and $\tilde{u}_{1}=\tilde{u}_{egr2}$. Here, $\tilde{u}_{1,0}=\tilde{u}_{egr1,0}$ and $\tilde{u}_{2,0}=\tilde{u}_{egr2,0}$ are the initial conditions and $u_{d}=u_{degr}=u_{egr}(t-\tau_{degr})$, where $\tau_{degr}$ is the time delay coefficient. The loss function for individual output variables is the mean square error between the true value ($\tilde{u}_{1}$) and the predicted one ($\tilde{u}_{1}(\bm{\theta})$), and the total loss is the weighted sum ($a_{1}=1$, $a_{2}=100$) of loss for both outputs. In the case of prediction of VGT actuator states, there are only two branch networks, branch 1 and 2 and only one output $\tilde{u}_{1}=\tilde{u}_{vgt}$. The inputs to these branch networks are $u_{d}=u_{dvgt}=u_{egr}(t-\tau_{dvgt})$ and the initial condition $\tilde{u}_{1,0}=\tilde{u}_{vgt,0}$. Appropriate scaling of the input and output are considered for fast and better training. Details of the network sizes and input-output data structures are shown in Table C.1.
Table 6: DeepONet results for independent variables: The training and testing errors for DeepONet for EGR and VGT states. The two DeepONets are trained using clean data and with added Gaussian noise. The relative $L_{2}$ error (%) for training and testing are shown in the fourth and sixth columns, respectively. The bottom part of the table gives the $L_{2}$ error (%) in training and testing when Gaussian noise is added to the training labelled data. We also note that the error in the case of noisy data is calculated with respect to the clean data. The mean and standard deviation of the error are calculated based on six different independent runs.
DeepONet Output Training Testing
Variable Data type Rel. $L_{2}$ error (%) Data type Rel. $L_{2}$ error (%)
$\mathcal{D}_{1}$ $\tilde{u}_{egr1}$ Clean data $0.64\pm 0.04$ Clean data $2.69\pm 0.49$
$\tilde{u}_{egr2}$ Clean data $0.65\pm 0.04$ Clean data $3.64\pm 0.18$
$\mathcal{D}_{2}$ $\tilde{u}_{vgt}$ Clean data $0.86\pm 0.10$ Clean data $2.52\pm 0.08$
with added noise in output during training
$\mathcal{D}_{1}$ $\tilde{u}_{egr1}$ 3% noise added $0.95\pm 0.08$ Clear data $3.11\pm 0.75$
$\tilde{u}_{egr2}$ 3% noise added $0.97\pm 0.02$ Clear data $3.48\pm 0.17$
$\mathcal{D}_{2}$ $\tilde{u}_{vgt}$ 3% noise added $1.14\pm 0.11$ Clear data $2.62\pm 0.13$
4 Results: PINN for inverse problem and transfer learning
In the previous sections, we discussed the results of pre-trained neural networks for empirical formulae and DeepONets for EGR and VGT states. We introduced a hybrid method that incorporates these pre-trained networks within the PINN framework in section 2.4 and hybrid multi-stage transfer learning in section 2.5. In this section, we present results for the inverse problem solved using the hybrid multi-stage transfer learning approach outlined in section 2.5. We consider two scenarios: i) training with clean data, and ii) training with noisy data. Since we use simulated data for model training, we added Gaussian noise to these simulated outputs to mimic real-world conditions. Based on expected noise from field sensors, we added Gaussian noise values of 3%, 3%, 1%, 5% and 3% to simulated outputs for $p_{im}$, $p_{em}$, $\omega_{t}$, $W_{egr}$ and $T_{em}$ respectively. The input signal trains required for the Simulink model viz. control input $u=\{u_{\delta},u_{egr},u_{vgt}\}$ and engine speed $n_{e}$, are obtained from field measurements on an engine test bed. As discussed in section 2.5, we train the hybrid PINN model on individual one minute data segments. In the first segment, the complete hybrid PINN model is trained. Complete training of this hybrid PINN model is computationally expensive since it requires learning all network parameters and requires more training epochs. To reduce computational time, we transfer the network weights and biases to the multi-stage transfer learning model for subsequent time segments. Refer Table 10 showing the improvement in computational performance with the transfer learning approach. For each minute of data segment, 301 residual points were considered in the physics-informed loss term of equation 21. Heuristically, we identified that with clean data, 151 residual points were sufficient for accurate parameter and state estimations. However, for noisy data, 301 residual points were used during the training process. The data sampling frequency was 5 Hz.
As discussed earlier in section 2.4, we approximate $p_{im}$, $p_{em}$, $\omega_{t}$, $T_{1}$ and $x_{r}$ using neural networks. We also incorporate mask function and scaling for the unknown parameters, as shown in Table 1. We note that output transformation, mask function, and scaling are important to ensure the network outputs and unknown parameters lie in $\mathbb{R^{+}}$. Additionally, we incorporate self-adaptive weights for the data loss term of $\bm{\lambda}_{p_{em}}$, $\bm{\lambda}_{p_{im}}$, $\bm{\lambda}_{\omega_{t}}$ and $\bm{\lambda}_{W_{egr}}$, initialized from a Gaussian random variable with mean $\dfrac{\alpha X_{data}}{\max(X_{data})}$, $X=\{p_{em},p_{im},\omega_{t},W_{egr}\}$ and standard deviation $\sigma=0.1$, where $\alpha$ is a heuristically determined weight multiplier ($10^{3}$, $10^{3}$, $10^{3}$, and $2\times 10^{3}$ for $\bm{\lambda}_{p_{em}}$, $\bm{\lambda}_{p_{im}}$, $\bm{\lambda}_{\omega_{t}}$ and $\bm{\lambda}_{W_{egr}}$ respectively).
The self-adaptive weight $\lambda_{T_{em}}$ and $\lambda_{T_{1}}$ are initialized from a Gaussian random variable with mean $1500$ and standard deviation 0.1. A softplus masking function is applied to the self-adaptive weights to ensure that they remain positive. The unknown parameters are initialized from a Gaussian random variable with standard deviation $\sigma=0.2$ and mean near the true values of the parameters. The parameters of the neural network, unknown engine parameters, and self-adaptive weights are optimized using the Adam optimizer [31] with piece-wise constant learning rates. To estimate the epistemic uncertainty, we train the hybrid PINN model separately for each data segment to generate an ensemble of 30 runs. This uncertainty arises from the random initialization of self-adaptive weights $\mathbf{\lambda}$, unknown parameter values $\mathbf{\Lambda}$, and dropouts included in the surrogate networks for estimating empirical formulae discussed in section 2.2. Furthermore, we transfer the network parameters from different source models (trained on segment $0-60$ sec) which also adds to the overall uncertainty. In the following sections, we present results for clean and noisy data conditions. A summary of the mean and standard deviation estimates for the unknown parameters under different test conditions is presented in table 7.
Table 7: Predicted unknown parameters (multi-stage Transfer Learning). The mean and standard deviation of the predicted unknown parameters when predicted using multi-stage transfer learning with clean data and Gaussian noisy data. The mask function is applied to the trainable variable to ensure positive values for predictions. Furthermore, scaling factors are used to convert engine parameter estimates from neural networks back to their original scales. For clean data, we show results obtained by using 151 and 301 data points while for noisy data, we show results obtained by using 301 data points. With noisy data, more data points are required for accurate parameter and state estimations.
Duration (sec) $A_{egrmax}$ $A_{vgtmax}$ $\eta_{sc}$ $h_{tot}$ Remarks
Parameter estimates, Clean data
$0\sim 60$ $1.009\pm 0.0017$ $1.022\pm 0.0003$ $1.004\pm 0.0008$ $1.003\pm 0.0035$ 151 data points for each known variables (for each 60 sec.).
$60\sim 120$ $0.922\pm 0.0487$ $1.0185\pm 0.0036$ $0.977\pm 0.0183$ $0.850\pm 0.0967$
$120\sim 180$ $0.955\pm 0.0117$ $0.991\pm 0.0036$ $1.012\pm 0.0017$ $1.013\pm 0.0041$
$180\sim 240$ $1.007\pm 0.0017$ $1.019\pm 0.0004$ $1.003\pm 0.0025$ $0.992\pm 0.0072$
$240\sim 300$ $0.944\pm 0.0980$ $0.999\pm 0.0073$ $1.001\pm 0.0025$ $0.988\pm 0.0103$
$0\sim 60$ $1.004\pm 0.0010$ $1.024\pm 0.0003$ $1.007\pm 0.000$ $1.005\pm 0.0012$ 301 data points for each known variables (for each 60 sec.).
$60\sim 120$ $0.971\pm 0.0470$ $1.012\pm 0.0046$ $1.000\pm 0.0833$ $0.980\pm 0.0561$
$120\sim 180$ $0.964\pm 0.0123$ $0.992\pm 0.0011$ $1.008\pm 0.0025$ $1.007\pm 0.0056$
$180\sim 240$ $1.004\pm 0.0013$ $1.019\pm 0.0003$ $1.003\pm 0.0017$ $0.992\pm 0.0054$
$240\sim 300$ $0.900\pm 0.1033$ $1.006\pm 0.0022$ $0.998\pm 0.0042$ $0.984\pm 0.0137$
Parameter estimates, Noisy data.
Noise level: $p_{em}=3\%$, $p_{im}=3\%$, $\omega_{t}=1\%$, $W_{egr}=5\%$, $T_{em}=3\%$
$0\sim 60$ $0.997\pm 0.1103$ $1.021\pm 0.0061$ $0.999\pm 0.0125$ $0.999\pm 0.0452$ 301 data points for each known variables (for each 60 sec.).
$60\sim 120$ $0.952\pm 0.0693$ $1.011\pm 0.0072$ $1.007\pm 0.0108$ $1.021\pm 0.0675$
$120\sim 180$ $0.803\pm 0.1047$ $0.974\pm 0.0148$ $0.998\pm 0.0125$ $0.984\pm 0.0734$
$180\sim 240$ $1.004\pm 0.0353$ $1.019\pm 0.0051$ $0.993\pm 0.0117$ $0.978\pm 0.0271$
$240\sim 300$ $0.793\pm 0.1333$ $1.002\pm 0.0103$ $0.997\pm 0.0117$ $0.985\pm 0.0328$
True value $1.000$ $1.000$ $1.000$ $1.000$ Normalized values for confidentiality
4.1 Parameter and state estimation under clean conditions
The estimates for the unknown parameters $\mathbf{\Lambda}$ obtained by training the hybrid PINN model for each data segment with the multi-stage transfer learning approach are illustrated in figure 6 and B.1.1. The variability in parameter estimation is obtained by computing the expected value and variance across thirty independent runs for each segment. The estimate of unknown parameters $\mathbf{\Lambda}$ represent the values obtained at the end of 35,000 epochs. We present a comparison between models trained with 151 and 301 residual points to validate our assertion that 151 points are adequate for accurate parameter estimation in clean data. Training with fewer data improves computational efficiency and hence, is preferred for field applications. From the violin plots in figure 6 and B.1.1, we note a high variation in all parameter estimates at t = 1-2 min. This is expected, as in this segment, both dynamic and idling conditions are encountered which pose challenges for engine state estimation using neural networks due to spectral bias [38]. With 301 data points, we observe a sightly reduced variance of parameter estimation across all $\mathbf{\Lambda}$ but with increased computation cost. This trade-off between requirements of computational time and accuracy needs to be considered for field application.
Figure 7 illustrates the prediction results for the eight output states defined in equations 3 through 10. Due to confidentiality, the dynamics shown in this figure are normalized on a scale 0-1 using the min-max values as:
$x_{\text{scale}}=\dfrac{x-x_{\text{min}}}{x_{\text{max}}-x_{\text{min}}}$ (26)
where $x_{\text{max}}$, $x_{\text{min}}$ maximum and minimum value of the true data of $x$ variable within the time duration considered. Here, it is noteworthy that states $\tilde{u}_{egr1},\tilde{u}_{egr2}$, and $\tilde{u}_{vgt}$ are predicted directly using the pre-trained DeepONet model. The PINN model prediction indicates good match with simulated ground truth data for states $p_{im},p_{em}$, and $\omega_{t}$. However, states $T_{1}$ and $x_{r}$, which are approximated by the PINN model, exhibit higher standard deviations from simulated ground truth data. This can be attributed to the relatively smaller influence of these two state variables on gas dynamics. Prediction for these two states depend solely on the gradient of the physics-based loss term which is small leading to higher variation in state estimation due to numerical error. Figure 8 shows the different gas flow rates obtained as intermediate states by using relationships between primary dynamics and other quantities of interest (QoI). Overall, the prediction from PINN model demonstrates good approximation to ground truth values for all states. We note that state $A_{egr}$ determines the area of EGR valve that is controlled by the EGR actuator. We observe that the prediction from PINN model for segments 1-2 min and 4-5 min show higher deviation as compared to true values. Since this state estimate is directly related to the estimate of the unknown parameter $A_{egrmax}$, we expect the parameter estimate to exhibit high variability in these segments as well. We observe this behavior in figure 6 and B.1.1, which depict higher variation for the parameter $A_{egrmax}$ in these segments. We show additional ensemble plots by combining all 30 runs across the five segments in appendix B section B.1. The error band in these ensemble plots help estimate the epistemic uncertainty from this transfer learning model. We note higher variation in prediction of states $T_{1}$ and $x_{r}$, that correlates with our discussion about these states earlier in this section. Figure 19(b) shows a convergence plot for the four unknown parameters of interest for this study along with the range of expected values for time segment 3-4 minute. The uncertainty band shown in this plot is obtained by calculating point wise minimum and maximum values of these parameters across thirty independent runs. Note that all the four parameters converge around the true value of these parameters across all runs, which demonstrates the repeatability of this hybrid transfer learning framework.
Fig. 6: Violin plot, multi-stage transfer learning, clean data, 151 data points showing the variation in unknown parameter predictions over 30 runs with clean data and 151 data points. 151 data points are considered for each of the known parameters (for each minute) under clean conditions. The violin plots are limited to the maximum and minimum values of the predicted parameters. The dots on the violin plots show a swarm plot indicating each predicted value for the 30 independent runs. The grey band shown indicates ± 7.5% error band. The dotted line in the centre indicates the true value of the unknown parameter. The top dotted line indicates 1.075 times the true value while the bottom dotted line indicates 0.925 times the true value.
Fig. 7: Predicted states, multi-stage transfer learning, clean data, 151 data points. Primary gas flow dynamics states from multi-stage transfer learning model (data scale normalized due to confidentiality). The plot shows a representative case where clean data is used with 151 data points for online training. Note that states $\tilde{u}_{egr1},\tilde{u}_{egr2},$ and $\tilde{u}_{vgt}$ are outputs obtained from pre-trained DeepONet and are not trained during the transfer learning process.
Fig. 8: Gas flow rates, multi-stage transfer learning, clean data, 151 data points. Gas flow rates from transfer learning model (data normalized due to confidentiality). The plot shows a representative case where clean data is used with 151 data points for online training. The gas flow rates are used for parameter estimation through flow equations defined in appendix A.
4.2 Parameter and state estimation under noisy conditions
To simulate a practical scenario, we assess the performance of our model when trained using noisy data. During our testing, we identified the need to use more data points to enhance the prediction accuracy of the transfer learning model. Consequently, we utilize 301 data points for training the PINN models under noisy conditions. Figure 9 illustrates the distribution of unknown parameters $\mathbf{\Lambda}$ over the five segments with thirty independent runs for each segment. With noise, we see higher variation across all parameter estimates which is expected due to the aleatoric uncertainty. Notably, amongst the four parameters, $A_{egrmax}$ demonstrates the largest increase in variation across all segments. In figure 10 and 11, we plot the prediction of primary gas dynamics and gas flow rates respectively for one sample run. We observe higher error in the prediction for state $A_{egr}$ when compared to its ground truth, correlating with the higher variability of $A_{egrmax}$ observed. Additional ensemble plots showing the predicted values and expected range of predictions for different QoIs are shown in Appendix B section B.2.
Fig. 9: Violin plot, multi-stage transfer learning, noisy data, 301 data points: Variation in engine parameter prediction over 30 runs with Gaussian noise in data signals with 301 data points. The dots on the violin plots shows a swarm plot indicating each predicted value for the 30 independent runs. The grey band shown indicates ± 7.5% error band. The dotted line in center indicates the true value of the unknown parameter. Higher variation in parameter estimation is observed across all time segments and represents the aleatoric uncertainty expected in field due to sensory noise.
Fig. 10: Predicted states, multi-stage transfer learning, noisy data, 301 data points: Prediction results for primary gas flow dynamics states from multi-stage transfer learning model (data scale normalized due to confidentiality). The plot shows a representative case where 301 noisy data points are used for online training of hybrid PINN model.
Fig. 11: Gas flow rates, multi-stage transfer learning, noisy data, 301 data points: Gas flow rates from multi-stage transfer learning model (data scale normalized due to confidentiality). The plot shows a representative case where 301 noisy data points are used for online training.
5 Results: Few-shot transfer learning
5.1 Training of multi-head neural networks
As discussed earlier, few-shot transfer learning involves two Phases: (1) offline training of the neural network with multiple heads and (2) training the head and unknown parameters for the inverse problem. The training in Phase I is done with 240 heads for each variable; each head is used for estimating a one-minute data segment. The input to the network is $0-60$ sec ($dt=0.2$ sec) normalized to $[-1,1]$. The first 120 heads correspond to the state variable for Case-I, and the second 120 heads correspond to the state variable for Case-IV (ref Table 5). We have shown the network detail and training error for each state variable in Table 8. We trained these networks using the Adam optimizer with an MSE loss function between the labelled and the predicted output. In the case of $N_{1}(t,\bm{\theta})$, as there are two outputs, an appropriate weight is considered while summing the two losses. The weighs of the network are initialized from a Gaussian distribution with mean zero and standard deviation $\sigma=\sqrt{2/(\text{dim}_{\text{in}}+\text{dim}_{\text{out}})}$ except for the last layer. The last layer is initialized with a standard deviation of $\sigma=\sqrt{10/(\text{dim}_{\text{in}}+\text{dim}_{\text{out}})}$. This shows a better accuracy compared to an initialization with a similar standard deviation as the hidden layers. This may be because, as $\text{dim}_{\text{out}}$ in the last layer has more neurons, the standard deviation becomes smaller compared to the previous layers. Furthermore, to check the applicability of the method for practical application, the training in Phase-I is done with different parameters and the transfer learning (Phase-II) is done with different parameters, as discussed in section 2.1 (ref Table 1).
Table 8: Training of multi-head networks for few-shot transfer learning (Phase-I): Multi-head neural networks considered to approximate the state variables and $x_{r}$ and $T_{1}$. The input to the network is time ($t$). The input time is 0 to 60 sec with $\delta t=0.2$ sec, and the input is scaled between $[-1,1]$. Each head approximates a minute of the variable. We considered 240 heads for each output approximating 240 minutes of dynamics. Thus, network $\mathcal{N}_{2}(t,\bm{\theta}_{2})$, $\mathcal{N}_{3}(t,\bm{\theta}_{3})$ and $\mathcal{N}_{4}(t,\bm{\theta}_{4})$ has 240 output, however network $\mathcal{N}_{1}(t,\bm{\theta}_{1})$ has 480 output as it approximate two ($p_{em}$ and $p_{im}$) variables. The ”Output transformation” and scaling are considered as per Table 1. The $L_{2}$ error shows % relative $L_{2}$ error in training for each variable.
Network $\mathcal{N}_{1}(t,\bm{\theta}_{1})$ $\mathcal{N}_{2}(t,\bm{\theta}_{2})$ $\mathcal{N}_{3}(t,\bm{\theta}_{3})$ $\mathcal{N}_{4}(t,\bm{\theta}_{4})$
Output $p_{em}$, $p_{im}$ $x_{r}$ $T_{1}$ $\omega_{t}$
[…]
5.2 Parameter and state estimation with clean data: Few shot method
After training,, the next step is to perform PINN to identify the engine parameters (Phase-II). The output layer ($\bm{W}_{out}$, $b_{out}$) of the multi-head networks is replaced with one output and trained for a new control vector ($\bm{u}=\{u_{\delta},u_{vgt},u_{egr}\}$) and engine speed ($n_{e}$). The weights of the output layer are initialized from a random Gaussian distribution with mean zero and standard deviation $\sigma=\sqrt{2/(\text{dim}_{\text{in}}+\text{dim}_{\text{out}})}$. The weights are sorted after initialization, and it shows better and faster convergence. The biases are initialized with zero. Similar to the PINN and transfer learning method, each of the self-adaptive weights for the data loss term of $\bm{\lambda}_{p_{em}}$, $\bm{\lambda}_{p_{im}}$, $\bm{\lambda}_{\omega_{t}}$ and $\bm{\lambda}_{W_{egr}}$ are initilized from a Gaussian random variable with mean $\dfrac{\alpha X_{data}}{\max(X_{data})}$, $X=\{p_{em},p_{im},\omega_{t},W_{egr}\}$ and standard deviation $\sigma=0.1$, where $\alpha$ is a initial weight given ($10^{3}$, $10^{3}$, $10^{3}$ and $2\times 10^{3}$ for $\bm{\lambda}_{p_{em}}$, $\bm{\lambda}_{p_{im}}$, $\bm{\lambda}_{\omega_{t}}$ and $\bm{\lambda}_{W_{egr}}$ respectively). The self-adaptive weight $\lambda_{T_{em}}$ and $\lambda_{T_{1}}$ are initialized from a Gaussian random variable with mean $20$ and standard deviation 0.1. We consider a softplus mask function to the self-adaptive weights so that the weights are always positive. The unknown parameters are initialized from a Gaussian random variable with standard deviation $\sigma=0.2$ and mean as $0.7$, $2.25$, $1.4$ and $2.5$ for $\eta_{sc}$, $h_{tot}$, $A_{egrmax}$ and $A_{vgtmax}$ respectively. Furthermore, as these quantities are always positive, we consider a mask function and scaling as shown in Table LABEL:Table:Mask_and_scaling_Unknown_parameter. Also note that the trainable parameters of the problem are the weights and bias of the last layer of the neural networks, the unknown parameters of the engine and the self-adaptive weights. These are optimized using eight different Adam optimizers: one for the weights and the biases of the neural networks, one for the unknown parameters, and one for each of the six self-adaptive weights. The loss function is given by the equation 20 as discussed in section 2. The learning rate scheduler is shown in Table LABEL:Table:Learniong_rate_schedule_few_shot (in Appendix C). We would also like to mention that we only optimized the self-adaptive weights up to $20\times 10^{3}$ epoch. After that, we fixed the self-adaptive weights with the value at $20\times 10^{3}$ epochs. The size of the self-adaptive weights for $\bm{\lambda}_{p_{em}}$, $\bm{\lambda}_{p_{im}}$, $\bm{\lambda}_{\omega_{t}}$ and $\bm{\lambda}_{W_{egr}}$ are same as that of the number of data point considered for each of them. Furthermore, as discussed earlier, in order to reduce the computational cost, we perform the derivative of the output with respect to the input outside the training loop as follows,
$y=S_{p}\left(\bm{H}(t)\bm{W}_{out}+b_{out}\right)$ (27a)
$\dfrac{dy}{dt}=\dfrac{1}{1+\exp[\bm{H}(t)\bm{W}_{out}+b_{out}]}\exp\left[\bm{H}(t)\bm{W}_{out}+b_{out}\right]\dfrac{d\bm{H}(t)}{dt}\bm{W}_{out}$ (27b)
where $S_{p}$ is softplus function (refer Table 8), $\bm{H}(t)$ is the output of the last hidden layer, $\bm{W}_{out}$ and $b_{out}$ are the weights and bias of the output layer, $\dfrac{d\bm{H}(t)}{dt}$ is the derivative of the output of the last hidden layer with respect to input $t$. From Eq. 27b, the derivative term is equal to $\dfrac{d\bm{H}(t)}{dt}$, and as the hidden layers are not trained, this remains fixed and can be evaluated analytically outside the network training loop. The size of $\dfrac{d\bm{H}(t)}{dt}$ is a matrix with size $n_{t}\times n_{h}$, where $n_{t}$ is the number of input points in time, and $n_{h}$ is the number of neurons in the last hidden layer.
We study two different scenarios with clean data without additional noise in the field data (simulated) for the inverse problem and a case with additional noise. We study five minutes of engine operation and sub-divide into five independent problems, each studied independently. For each one minute, we consider 301 residual points with equal $dt=0.2$ sec. For the first case of clean data, we consider $151$ data points ($dt=0.4$ sec.) for the data loss of $p_{em}$, $p_{em}$, $\omega_{t}$, $W_{egr}$ and $T_{em}$. As discussed earlier, the size of self-adaptive weight in this case is $151\times 1$ for $\lambda_{p_{em}}$, $\lambda_{p_{im}}$, $\lambda_{\omega_{t}}$ and $\lambda_{W_{egr}}$, the size self-adaptive weight for $\lambda_{T_{em}}$ and $\lambda_{T_{1}}$ is $1\times 1$. For clean data, we consider $301$ data points ($dt=0.2$ sec) for calculating data loss for $p_{em}$, $p_{em}$, $\omega_{t}$, $W_{egr}$ and $T_{em}$. The size of the self-adaptive weights is $301$ for $\lambda_{p_{em}}$, $\lambda_{p_{im}}$, $\lambda_{\omega_{t}}$ and $\lambda_{W_{egr}}$ while the size of self-adaptive weight for $\lambda_{T_{em}}$ and $\lambda_{T_{1}}$ is $1\times 1$. We note that despite using clean data, uncertainty is included in our model due to use of dropouts in pre-trained networks. Furthermore, in order to check the effect of initialization on the accuracy, we consider $30$ independent runs, for each one minute data segment.
Table 9 shows the mean and standard deviation of the predicted unknowns for each time segment. The corresponding violin plots are shown in Figs. 12 and B.3.1 for clean data with 151 and 301 points points respectively. One of the realizations of the dynamics of the predicted gas flow states and $x_{r}$ and $T_{1}$ with 151 data points is shown in Fig. 13. Similarly, we show the corresponding intermediate state variables in Fig. 14. All these results are normalized to a scale 0-1 using Eq. 26 for confidentiality. From Table 9 and Figs. 12 and B.3.1, it can be observed that predicted parameters show a good accuracy except during the 2-3 minutes duration. As discussed earlier, this may be due to an engine idle case during this time. We also observe outliers in $h_{tot}$ prediction during 1-2 minute when 151 data points are considered, while this is reduced when we considered 301 data points.
Fig. 12: Violin plot, few-shot learning, clean data, 151 data points: of the predicted value of the unknown parameters. 151 data points are considered for each of the known parameters (for each one minute segment) and no noise is considered in this case. The violin plots are clipped at the maximum and minimum values of the predicted parameters. The results are shown normalizing the true value of the unknown parameters to be one. The dots over the violin plots show the swarm plot indicates the predicted value for individual runs. The grey region shown between the dotted lines indicates the $\pm$ 7.5% error band. The middle dotted line indicates the true value of the unknown parameter, the top dotted line indicates 1.075 times the true value while the bottom dotted line indicates 0.925 times the true value.
Table 9: Predicted unknown parameters (Few-shot learning): The mean and standard deviation of the predicted unknown parameters when predicted using Few-shot transfer learning with clean data and Gaussian noisy data. The first set of results are predicted mean $\pm$ standard deviation when 151 data points for the known variables (for each 60 sec segment) with clean data, the second set of results are when 301 data points for the known variables (for each 60 sec). The third set of results for noisy data with 301 data points for each known variables (for each 60 sec segment).
Duration (sec) $A_{egrmax}\times 10^{-4}$ $A_{vgtmax}\times 10^{-4}$ $\eta_{sc}$ $h_{tot}$ Remarks
Parameter estimates, Clean data
$0\sim 60$ $1.0073\pm 0.0046$ $1.0201\pm 0.0008$ $0.9929\pm 0.0086$ $0.9592\pm 0.0296$ Clean data, 151 data points for each known variables (for each 60 sec.).
$60\sim 120$ $0.9531\pm 0.0026$ $1.0115\pm 0.0007$ $0.9980\pm 0.0162$ $1.0559\pm 0.1541$
$120\sim 180$ $0.9512\pm 0.0238$ $0.9925\pm 0.0022$ $0.9281\pm 0.0357$ $0.7854\pm 0.1144$
$180\sim 240$ $1.0057\pm 0.0013$ $1.0186\pm 0.0006$ $0.9755\pm 0.0091$ $0.9589\pm 0.0173$
$240\sim 300$ $0.9238\pm 0.0032$ $0.9993\pm 0.0018$ $0.9965\pm 0.0043$ $0.9951\pm 0.0163$
$0\sim 60$ $1.0038\pm 0.0044$ $1.0198\pm 0.0007$ $0.9922\pm 0.0056$ $0.9599\pm 0.0207$ Clean data, 301 data points for each known variables (for each 60 sec.).
$60\sim 120$ $1.0312\pm 0.0042$ $1.0043\pm 0.0008$ $0.9958\pm 0.0111$ $1.0084\pm 0.0883$
$120\sim 180$ $0.9551\pm 0.0263$ $0.9922\pm 0.0014$ $0.9385\pm 0.0315$ $0.8205\pm 0.1001$
$180\sim 240$ $1.0057\pm 0.0013$ $1.0190\pm 0.0006$ $0.9688\pm 0.0098$ $0.9532\pm 0.0162$
$240\sim 300$ $0.9385\pm 0.0041$ $1.0024\pm 0.0006$ $0.9893\pm 0.0047$ $0.9646\pm 0.0156$
Predicted unknown, Noisy data.
Noise level: $p_{em}=3\%$, $p_{im}=3\%$, $\omega_{t}=1\%$, $W_{egr}=5\%$, $T_{em}=3\%$
$0\sim 60$ $0.9699\pm 0.0653$ $1.0190\pm 0.0056$ $0.9913\pm 0.0094$ $0.9568\pm 0.0379$ Noisy data, 301 data points for each known variables (for each 60 sec.).
$60\sim 120$ $1.0255\pm 0.0748$ $0.9992\pm 0.0092$ $0.9888\pm 0.0230$ $0.9445\pm 0.1749$
$120\sim 180$ $0.5800\pm 0.0568$ $0.9520\pm 0.0165$ $0.9524\pm 0.0429$ $0.9087\pm 0.2365$
$180\sim 240$ $0.9972\pm 0.0246$ $1.0139\pm 0.0081$ $0.9554\pm 0.0276$ $0.9114\pm 0.0721$
$240\sim 300$ $0.9525\pm 0.0998$ $0.9942\pm 0.0093$ $0.9849\pm 0.0122$ $0.9519\pm 0.0390$
True value $1.000$ $1.000$ $1.000$ $1.000$
From Fig. 13, it can be observed that neural network predictions show a good agreement with the true value except $x_{r}$ and $T_{1}$. Similar to the multi-stage transfer learning cases, we observed that the effect of $x_{r}$ and $T_{1}$ on the derived variable are minimum, as can be seen from the results of $T_{e}$ and $T_{em}$ in Fig. 14. We have shown the different intermediate states by using the relationships between the primary state variables (NN output) and predicted known and constants in Fig. 14. These variables also show good agreement with the true value. As discussed earlier, we have run the model 30 times for each time zone, and the ensemble data are shown in Appendix B.3. The spikes visible in the predicted results are due to the dropout introduced in the pre-trained networks.
5.3 Parameter and state estimation with noisy data: Few shot method
In the previous section, we discussed the results for the predicted parameters and the state variable with clean data for the field variables. i.e. we did not consider any added noise to the simulated data. However, field data may be noisy in practice due to the inherent instrument noise or other reasons. Thus, to simulate a more practical scenario and the applicability of the proposed method, we consider added noise and perform PINN. We added a Gaussian noise of 3%, 3%, 1%, 5% and 3% added to the simulated data of $p_{em}$, $p_{im}$, $\omega_{t}$, $W_{egr}$ and $T_{em}$ respectively. In this case, we consider 301 data points. The other parameters are kept the same as discussed in the previous section for clean data. Before discussing further, we would like to mention that in this case, there are three different types of uncertainties: (i) uncertainty associated with the initialization of the last layer of the neural network and other parameters (e.g. unknown parameters, self-adaptive weights), (ii) uncertainty in the pre-trained network due to the dropout considered, (iii) uncertainty due to the added noise in the field variables. The mean $\pm$ standard deviation of the predicted unknown parameters for each time segment is shown in Table 9 and the corresponding violin plots are shown in Fig. 15. We observed that the standard deviation of the predicted unknown has increased compared to the clean data. This is expected due to aleatory uncertainty due to additional noise in data. Similar to the clean data case, the results in the 2$\sim$3 minute duration show a higher deviation due to engine idling scenario.
Fig. 13: Predicted states, Few-shot learning, Clean data, 151 data points: One of the predicted realization of the state variables and $x_{r}$ and $T_{1}$, for the case where clean data and 151 data points are considered for each known variables (for each one minute duration). Each 1 minute is trained independently and combined to get the total 5 minutes results. The vertical dotted blue lines separate each minute result. The results are normalized using Eq. 26 with maximum and minimum true value calculated over the 5 minute.
Fig. 14: Other variables, Few-shot learning, Clean data, 151 data points: One of the predicted realization of different variables of the diesel engine for the case where clean data and 151 data points are considered for each known variable (for each one minute duration). The variable $A_{egr}$, $T_{e}$, $T_{em}$ and $W_{t}$ are the variable which are directly depends on the unknown parameters $A_{egrmax}$, $\eta_{sc}$, $h_{tot}$ and $A_{vgtmax}$. The variable $T_{e}$ also depends on the variable $T_{1}$ and $x_{r}$.
Fig. 15: Violin plot, few-shot learning, noisy data, 301 data points: of the predicted value of the unknown parameters. 151 data points are considered for each of the known parameters (for each 1 minute) and no noise is considered. The violin plots are cut at the maximum and minimum values of the predicted parameters. The results are shown normalizing the true value of the unknown parameters to be one. The dots over the violin plots show the swarm plot indicating each predicted value. The grey colour shedding shown between the dotted lines indicates the $\pm$ 7.5% error band. The middle dotted line indicates the true value of the unknown parameter, the top dotted line indicates 1.075 times the true value while the bottom dotted line indicates 0.925 times the true value.
Fig. 16: Predicted states, Few-shot learning, Noisy data, 301 data points. One of the predicted realization of the state variables and $x_{r}$ and $T_{1}$, for the case where clean data and 151 data points are considered for each known variables (for each one minute duration). Each 5 minute are trained independently and combined to get the total 5 minute results. The vertical dotted blue lines separate each results. The results are normalized using Eq. 26 with maximum and minimum true value calculated over the 5 minute.
We study the dynamics of the state variables when we add noise to the field variables. As discussed earlier, we would also like to mention again that we did consider any noise in the input data (i.e. input control vector $\bm{u}=\{u_{egr},u_{vgt},u_{delta}\}$ and engine speed $n_{e}$) as the input data are collected from actual engines running condition. The dynamics of one of the realizations of the predicted state variables are shown in Fig. 16. It can be observed that even with noisy data, the predicted state variables show a good agreement with the true value. We also study the intermediate state variables, and the corresponding intermediate variables are shown in Fig. 17. It is observed agreement in the predicted results except in $A_{egr}$ in 2$\sim$3 time zone. However, the dynamic pattern matches even in this time zone. This is because in the time zone, the predicted $A_{egrmax}$ deviates from the true value for this particular realization. The little noise observed in the predicted results is excepted as we are considering noisy data. The predicted empirical formulae using the pre-trained neural networks are show in Fig. B.4.3. Similar to the clean data case, we observed noise in the predicted empirical formulae due to dropout. However, in this case, the noise is more as we have considered added noise in the field data. We have shown the ensemble data in the Appendix.
Fig. 17: Other variables, Few-shot learning, Noisy data, 301 data points. One of the predicted realization of different variables of the diesel engine for the case where clean data and 151 data points are considered for each known variable (for each one minute duration). The variable $A_{egr}$, $T_{e}$, $T_{em}$ and $W_{t}$ are the variable which are directly depends on the unknown parameters $A_{egrmax}$, $\eta_{sc}$, $h_{tot}$ and $A_{vgtmax}$. The variable $T_{e}$ also depends on the variable $T_{1}$ and $x_{r}$.
6 Comparison between Multi-stage transfer learning and Few-shot transfer learning
We have presented two approaches for efficiently identifying unknown parameters of a diesel engine using a hybrid PINNs based approach. In this section, we compare the two methods based on empirical evidence from our numerical experiments. We note that both methods yield comparable results in terms of the mean values of unknown parameters. However, the multi-stage transfer learning approach shows a higher variation in parameter estimation in general as compared to Few-shot transfer learning. This is attributed to the fact that the few shot transfer learning network was trained on a much larger dataset that captures a wide variety of engine operating conditions. Consequently, generalization is better with the few-shot approach as compared to the multi-stage transfer learning approach. For noisy data, we observe a much wider spread in the parameter estimation from the multi-stage transfer learning approach as compared to few-shot learning. This could be indicative of the need to retrain more layers of the transfer learning network instead of the last output layer to improve system dynamics approximation. For a particular time segment of $2-3$ minute, we observe that the estimation for $A_{egrmax}$ from the few-shot learning paradigm is less accurate as compared to the transfer learning approach. This might be because the flow through the EGR ($W_{egr}$) is very small (negligible) compared to the other time zone. Additionally, we do not observe a significant time difference in training the network with 151 or 301 data points. Any variations in computation time are transients due to network latency and the existing state of the computing node. Table 10 shows the computational time comparison between the baseline PINN model (hybrid PINN+DeepONet), multi-stage transfer learning, and few-shot learning paradigms. We note a significant improvement in computational time for the hybrid PINN methods we presented earlier compared to the standard/baseline PINN method presented in [20]. Fig. 18 shows the loss convergence for the ODE and data loss terms for different engine states and gas flows. We note that all loss terms for the two transfer learning paradigms converge earlier than the baseline PINN model to fully converge. Additionally, Fig. 19 shows the convergence plot for the four unknown engine parameters generated through an ensemble across thirty independent runs for time segment $t=3-4$ min. All unknown parameters for the transfer learning methods converge at $\approx 20K$ epochs, while for the baseline PINN model, this takes $\approx 50K$ epochs. Also, compared to the multi-stage transfer learning approach, the few-shot transfer learning method uses significantly lower computational time. This is because the few-shot method eliminates the need for automatic differentiation for the output variable. However, this approach requires a larger corpus of labeled data for the baseline training of hidden layers (body) of the PINN networks, which are trained offline. Hence, trade-offs between data availability, computational and accuracy requirements are essential for choosing one method over the other.
(a) Baseline PINN
(b) Multi-stage transfer learning
(c) Few shot learning
Fig. 18: Convergence of physics loss and data loss vs epoch for different method: (a) Baseline PINN (hybrid PINN) model, (b) Multi-stage transfer learning model, and (c) Few-shot transfer learning model. The losses plotted are without the self-adaptive weights. Losses for both the transfer learning methods converges earlier than the baseline PINN. Initial condition losses are not shown.
Table 10: Computational time comparison, multi-stage transfer learning vs few-shot. Mean ($\mu$) and standard deviation ($\sigma$) of computational time are calculated from 30 independent runs for each time segment. We see a significant improvement by using the multi-stage transfer learning approach when compared against the baseline PINN (hybrid PINN) model (which we train from $0-1$ min time segment). While the hybrid multi-stage transfer learning approach reduces computational time by approximately 50%, the few-shot learning paradigm further improves the computational time by an additional 25% in comparison to the baseline PINN model. Random variation in computational time results from system level variations and latency that exist in high-performance cluster systems which were used for this study. All computations were performed with 32-bit data precision on one CPU core with Tensorflow version 1.
Method Training epochs Training segment Time in sec ($\mu$, $\sigma$)
Clean data Noisy data
151 data points 301 data points
Baseline PINN (hybrid) 80,000 0-60s (487.12, 54.35) (585.99, 79.87)
Multi-stage Transfer learning 35,000 60-120s (245.50, 25.82) (235.08, 42.45)
120-180s (257.16, 18.49) (222.45, 51.76)
180-240s (219.47, 34.63) (242.05, 31.99)
240-300s (237.10, 26.26) (219.97, 45.19)
Few-shot Transfer learning 30,000 0-60s (124.35, 11.00) (140.28, 16.95)
60-120s (129.21, 18.20) (137.13, 11.72)
120-180s (128.48, 14.63) (133.41, 16.87)
180-240s (122.82, 12.28) (141.81, 13.31)
240-300s (125.71, 11.74) (135.28, 19.45)
7 Summary and future work
In this study, we developed two computationally efficient and practical PINN-based methods for identifying unknown parameters and dynamics of diesel engines. We optimized the problem at multiple levels, first introducing systematic (e.g. manufacturing) variability by adding dropout in the pre-trained networks used for estimating empirical quantities of the mean value gas flow model. Secondly, we proposed a hybrid PINN and operator-based (PINNs and DeepONet) framework as a step towards meeting real time computation requirements for field deployment. To reduce the computational time further, we introduced two transfer learning approaches: i) multi-stage transfer learning based on transferring knowledge from one section of signal to the remaining sections and ii) few-shot transfer learning where one network is trained offline on a large dataset comprising various operating conditions, and thereafter, knowledge is transferred for predicting the gas flow states for test signals. Both methods are suited for identifying unknown parameters using the known gas dynamics and appropriate sensor data, and we demonstrate this by using simulated data for different operating conditions. However, the input data to generate these simulated data are considered from actual engine running conditions. We also assess the sensitivity of our methods to field noise by adding Gaussian noise to our simulated data. Both methods perform reasonably well under both noise-free and noisy conditions. However, under noisy conditions, the estimates of unknown parameters and dynamic states show higher variability as compared to noise-free operating conditions. In terms of computational cost, the multi-stage transfer learning approach requires fine tuning on each segment of signal and is more expensive as compared to fine tuning the few-shot model that shares knowledge with a model trained on a larger data corpus. However, a higher computational cost is incurred on the training of the body (hidden layers) for the few-shot model upfront which is done offline.
One source of epistemic uncertainty in our model is the use of dropout rates in pre-trained empirical networks. In this study, we use dropout across our experiments. Understanding the sensitivity of our model’s predictions to changes in these dropout values will offer further insights into the epistemic uncertainty expected from our methods. The dropout rate should be estimated from uncertainty in the actual data so that the network prediction gives similar uncertainty. Furthermore, we also note that since dropout rate depends on network size, further study is necessary to determine the appropriate dropout values with network size. Additionally, methods to reduce the variability that we see in parameter estimation under noisy conditions is a necessary step for improving the performance of our methods when deployed on field units. Minimizing the effect of aleatoric uncertainty on the model’s prediction is critical to successful implementation of such methods in the field and we continue to investigate means for this. Another future direction may include the study of a computationally effective method for dynamical stoping of training (number of training epochs) for different operating conditions.
(a) Baseline PINN (hybrid PINN)
(b) Multi-stage transfer learning
(c) Few-shot transfer learning
Fig. 19: Parameter convergence, clean data, 151 data points, ensemble: Convergence of the four unknown parameters of the engine showing the range of variation across the 30 independent runs for (a) Baseline PINN (hybrid PINN) model ($t=0-1$ min), (b) Multi-stage transfer learning model ($t=3-4$ min), and (c) Few-shot transfer learning model ($t=3-4$ min) for clean data case with 151 data points in the loss function. The min-max band indicates the range over which the parameter estimation varies across the thirty independent runs that we used in this study. We observed that this variation is reduced significantly as training progresses, indicating a good repeatability of our PINN framework.
Input data and data generations
The input data are collected from actual engine running conditions. These data are used to generate simulated data with different ambient conditions using the Simulink file accompanied [21] and available at [24].
Author Contributions Statement
Kamaljyoti Nath: Conceptualization, Methodology, Software, Validation, Investigation, Visualization, Result generation and analysis, Writing- original draft; Varun Kumar: Validation, Investigation, Visualization, Result generation and analysis, Writing- original draft; Daniel J. Smith: Conceptualization, Data curation, Project administration, Supervision, Writing - original draft; George Em Karniadakis: Conceptualization, Methodology, Funding acquisition, Project administration, Resources, Supervision, Writing - original draft. All authors reviewed the manuscript.
Acknowledgement
This research was conducted using computational resources and services at the Center for Computation and Visualization, Brown University. This study was funded by Cummins Inc. USA.
Conflict of interest
Author George Em Karniadakis has received research support from Cummins Inc. Karniadakis has financial interests with the company PredictiveIQ. The rest of the authors declare no competing interests.
Appendices
Appendix A Engine model
In this section we present the complete numerical model for the mean value model considered in this study. We have considered the mean value engine model proposed by Wahlström and Eriksson [21] in our present study. We only present the equation here, more detail can be found in ref [21].
Manifold pressures
$\dfrac{d}{dt}p_{im}=\dfrac{R_{a}T_{im}}{V_{im}}(W_{c}+W_{egr}-W_{ei})$ (A.1)
$\dfrac{d}{dt}p_{em}=\dfrac{R_{e}T_{em}}{V_{em}}(W_{eo}-W_{t}-W_{egr})$ (A.2)
Cylinder
$W_{ei}$ $=\dfrac{\eta_{vol}p_{im}n_{e}V_{d}}{120R_{a}T_{im}}$ (A.3)
$W_{eo}$ $=W_{f}+W_{ei}$ (A.4)
$W_{f}=\dfrac{10^{-6}}{120}u_{\delta}n_{e}n_{cyl}$ (A.5)
$\eta_{vol}=c_{vol1}\sqrt{p_{im}}+c_{vol2}\sqrt{n_{e}}+c_{vol3}$ (A.6)
$T_{e}=\eta_{sc}\Pi_{e}^{1-1/\gamma_{a}}r_{c}^{1-\gamma_{a}}x_{p}^{1/\gamma_{a}-1}\left[q_{in}\left(\dfrac{1-x_{cv}}{c_{pa}}+\dfrac{x_{cv}}{c_{Va}}\right)+T_{1}r_{c}^{\gamma_{a}-1}\right]$ (A.7)
$\Pi_{e}=\dfrac{p_{em}}{p_{im}}$ (A.8)
$T_{1}=x_{r}T_{e}+(1-x_{r})T_{im}$ (A.9)
$x_{r}=\dfrac{\Pi_{e}^{1/\gamma_{a}}x_{p}^{-1/\gamma_{a}}}{r_{c}x_{v}}$ (A.10)
$x_{p}=$ $\dfrac{p_{3}}{p_{2}}=1+\dfrac{q_{in}x_{cv}}{c_{Va}T_{1}r_{c}^{\gamma_{a}-1}}$ (A.11)
$x_{v}=$ $\dfrac{v_{3}}{v_{2}}=1+\dfrac{q_{in}(1-x_{cv})}{c_{pa}[(q_{in}x_{cv}/c_{va})+T_{1}r_{c}^{\gamma_{a}-1}]}$ (A.12)
$q_{in}=\dfrac{W_{f}q_{HV}}{W_{ei}+Wf}(1-x_{r})$ (A.13)
$x_{r}=\dfrac{\Pi_{e}^{1/\gamma_{a}}x_{p}^{-1/\gamma_{a}}}{r_{c}x_{v}}$ (A.14)
$T_{em}=T_{amb}+(T_{e}-T_{amb})\exp\left(\dfrac{-h_{tot}\pi d_{pipe}l_{pipe}n_{pipe}}{W_{eo}c_{pe}}\right)$ (A.15)
EGR valve
$\dfrac{d\tilde{u}_{egr1}}{dt}$ $=\dfrac{1}{\tau_{egr1}}\left[u_{egr}(t-\tau_{degr})-\tilde{u}_{egr1}\right]$ (A.16)
$\dfrac{d\tilde{u}_{egr2}}{dt}$ $=\dfrac{1}{\tau_{egr2}}\left[u_{egr}(t-\tau_{degr})-\tilde{u}_{egr2}\right]$ (A.17)
$\tilde{u}_{egr}$ $=K_{egr}\tilde{u}_{egr1}-(K_{egr}-1)\tilde{u}_{egr2}$ (A.18)
$W_{egr}=\dfrac{A_{egr}p_{em}\Psi_{egr}}{\sqrt{T_{em}R_{e}}}$ (A.19)
$\Psi_{egr}=1-\left(\dfrac{1-\Pi_{egr}}{1-\Pi_{egropt}}-1\right)^{2}$ (A.20)
$A_{egr}=A_{egrmax}f_{egr}(\tilde{u}_{egr})$ (A.21)
$\Pi_{egr}=\begin{cases}\Pi_{egropt}\;\;\;\;\;\;&\text{if}\;\;\dfrac{p_{im}}{p_{em}}<\Pi_{egropt}\\ \dfrac{p_{im}}{p_{em}}\;\;\;\;\;\;&\text{if}\;\;\Pi_{egropt}\leq\dfrac{p_{im}}{p_{em}}\leq 1\\ 1\;\;\;\;\;\;&\text{if}\;\;1<\dfrac{p_{im}}{p_{em}}\end{cases}$ (A.22)
$f_{egr}(\tilde{u}_{egr})=\begin{cases}c_{egr1}\tilde{u}_{egr}^{2}+c_{egr2}\tilde{u}_{egr}+c_{egr3}\;\;\;&\text{if}\;\;\tilde{u}_{egr}\leq-\dfrac{c_{egr2}}{2c_{egr1}}\\ c_{egr3}-\dfrac{c_{egr2}^{2}}{4c_{egr1}}\;\;\;&\text{if}\;\;\tilde{u}_{egr}>-\dfrac{c_{egr2}}{2c_{egr1}}\end{cases}$ (A.23)
Turbocharger
$\dfrac{d}{dt}\omega_{t}=\dfrac{P_{t}\eta_{m}-P_{c}}{J_{t}\omega_{t}}$ (A.24)
$\dfrac{d\tilde{u}_{vgt}}{dt}=\dfrac{1}{\tau_{vgt}}\left(u_{vgt}(t-\tau_{dvgt})-\tilde{u}_{vgt}\right)$ (A.25)
$W_{t}=\dfrac{A_{vgtmax}p_{em}f_{\Pi_{t}}(\Pi_{t})f_{vgt}(\tilde{u}_{vgt}))}{\sqrt{T_{em}R_{e}}}$ (A.26)
$f_{\Pi_{t}}(\Pi_{t})=\sqrt{1-\Pi_{t}^{K_{t}}}$ (A.27)
$\Pi_{t}=\dfrac{p_{amb}}{p_{em}}$ (A.28)
$\left[\dfrac{f_{vgt}(\tilde{u}_{vgt})-c_{f2}}{c_{f1}}\right]^{2}+\left[\dfrac{\tilde{u}_{vgt}-c_{vgt2}}{c_{vgt1}}\right]^{2}=1$ (A.29)
$f_{vgt}(\tilde{u}_{vgt})=c_{f2}+c_{f1}\sqrt{\text{max}\left(0,1-\left(\dfrac{\tilde{u}_{vgt}-c_{vgt2}}{c_{vgt1}}\right)^{2}\right)}$ (A.30)
Here we define a variable $F_{vgt,\Pi_{t}}$ as
$F_{vgt,\Pi_{t}}=f_{vgt}(\tilde{u}_{vgt})\times f_{\Pi_{t}}(\Pi_{t})$ (A.31)
$P_{t}\eta_{m}=\eta_{tm}W_{t}c_{pe}T_{em}\left(1-\Pi_{t}^{1-1/\gamma_{e}}\right)$ (A.32)
$\eta_{tm}=\eta_{tm,max}-c_{m}(BSR-BSR_{opt})^{2}$ (A.33)
$BSR=\dfrac{R_{t}\omega_{t}}{\sqrt{2c_{pe}T_{em}(1-\Pi_{t}^{1-1/\gamma_{e}})}}$ (A.34)
$c_{m}=c_{m1}[max(0,\omega_{t}-c_{m2}]^{c_{m3}}$ (A.35)
Compressor
$\eta_{c}=\dfrac{P_{c,s}}{P_{c}}=\dfrac{T_{amb}(\Pi_{c}^{1-1/\gamma_{a}}-1)}{T_{c}-T_{amb}}$ (A.36)
$\Pi_{c}=\dfrac{p_{im}}{p_{amb}}$ (A.37)
$P_{c,s}=W_{c}c_{pa}T_{amb}(\Pi_{c}^{1-1/\gamma_{a}}-1)$ (A.38)
$P_{c}=\dfrac{P_{c,s}}{\eta_{c}}=\dfrac{W_{c}c_{pa}T_{amb}}{\eta_{c}}(\Pi_{c}^{1-1/\gamma_{a}-1})$ (A.39)
$\eta_{c}=\eta_{cmax}-\bm{\mathcal{X}}^{T}\bm{Q}_{c}\bm{\mathcal{X}}$ (A.40)
$\bm{\mathcal{X}}=\begin{bmatrix}W_{c}-W_{copt}\\ \pi_{c}-\pi_{copt}\end{bmatrix}$ (A.41)
$\pi_{c}=(\Pi_{c}-1)^{c_{\pi}}$ (A.42)
$\bm{Q}_{c}=\begin{bmatrix}a_{1}&a_{3}\\ a_{3}&a_{2}\end{bmatrix}$ (A.43)
$\Psi_{c}=\dfrac{2c_{pa}T_{amb}(\Pi_{c}^{1-1/\gamma_{a}}-1)}{R_{c}^{2}\omega_{t}^{2}}$ (A.44)
$\Phi_{c}=\dfrac{W_{c}/\rho_{amb}}{\pi R_{c}^{3}\omega_{t}}=\dfrac{R_{a}T_{amb}}{p_{amb}\pi R_{c}^{3}\omega_{t}}W_{c}$ (A.45)
$c_{\Psi 1}(\omega_{t})(\Psi_{c}-c_{\Psi 2})^{2}+c_{\Phi 1}(\omega_{t})(\Phi_{c}-c_{\Phi 2})^{2}=1$ (A.46)
$c_{\Psi 1}(\omega_{t})=c_{\omega_{\Psi 1}}\omega_{t}^{2}+c_{\omega_{\Psi 2}}\omega_{t}+c_{\omega_{\Psi 3}}$ (A.47)
$c_{\Phi 1}(\omega_{t})=c_{\omega_{\Phi 1}}\omega_{t}^{2}+c_{\omega_{\Phi 2}}\omega_{t}+c_{\omega_{\Phi 3}}$ (A.48)
$W_{c}=\dfrac{p_{amb}\pi R_{c}^{3}\omega_{t}}{R_{a}T_{amb}}\Phi_{c}$ (A.49)
$\Phi_{c}=\sqrt{\text{max}\left(0,\dfrac{1-c_{\psi 1}(\Psi_{c}-c_{\Psi 2})^{2}}{c_{\Phi 1}}\right)}+c_{\Phi 2}$ (A.50)
For the value considered for the parameters in the equation and general workflow, interested readers may refer to our previous work [20].
Appendix B Additional figures and Table
B.1 Multi-stage transfer learning prediction results for clean data
We provide additional results showing the predictions for different engine related QOIs. The plots combine the results from 30 independent runs and represent an estimate of the epistemic uncertainty of the network.
Fig. B.1.1: Violin plot, multi-stage transfer learning, clean data, 301 data points showing the variation in engine parameter predictions over 30 runs with clean data and 301 data points. More data points help reduce the variation in estimation across all data segments as compared to the model trained with 151 data points.
Fig. B.1.2: Predicted dynamic states, multi-stage transfer learning, clean data, 151 data points, ensemble: Ensemble prediction results for primary gas flow dynamics states from transfer learning model (data scale normalized due to confidentiality). The plot shown is for scenario where clean data is used with 151 data points used for online training. The ensemble mean (red) for the 30 runs is plotted against ground truth (black) for the eight dynamic states defined by ODE’s describing the gas flow dynamics. Error bands (shown in light green) indicate the spread of data between point-wise maximum and minimum values across all 30 runs.
Fig. B.1.3: Predicted gas flow, multi-stage transfer learning, clean data, 151 data points, ensemble: Ensemble prediction results for gas flow rates from the Transfer learning model with clean data over 30 independent runs across all segments. 151 data points were used during online parameter estimation task.
Fig. B.1.4: Predicted empirical states, multi-stage transfer learning, clean data, 151 data points, ensemble: Ensemble prediction estimates for empirical relationships for clean data with Transfer learning model. Note that in region $\approx 1.5-3.5\,min$, the model’s prediction indicates a spectral bias in learning the high frequency signals that typically result from engine idling. Elsewhere, the model predictions show good match with target values for these empirical relationships.
B.2 Multi-stage transfer learning prediction results for noisy data
Fig. B.2.1: Predicted dynamic states, multi-stage transfer learning, noisy data, 301 data points, ensemble: Ensemble prediction results for primary gas flow dynamics states from Transfer learning model (data scale normalized due to confidentiality). The plot shown is for scenario where noisy data is used with 301 data points used for online training. The ensemble mean (red) for the 30 runs is plotted against ground truth (black) for the eight dynamic states defined by ODE’s describing the gas flow dynamics. Error bands (shown in light green) indicate the spread of data between point-wise maximum and minimum values across all the 30 runs.
Fig. B.2.2: Predicted gas flow rate, multi-stage transfer learning, noisy data, 301 data points, ensemble: Ensemble prediction results for gas flow rates from the Transfer learning model with noisy data. 301 data points were used during online parameter estimation task.
Fig. B.2.3: Predicted empirical states, multi-stage transfer learning, noisy data, 301 data points, ensemble: Ensemble prediction estimates for empirical relationships for noisy data with transfer learning model with 301 data points. Note that in region $\approx 1.5-3.5\,min$, the model’s prediction indicate a spectral bias in learning the high frequency signals that typically result from engine idling. Also, there is a significant increase in expected variability in prediction of these empirical states under noisy conditions as compared to clean conditions shown in figure B.1.4, correlating with increase in variability in prediction of other states and parameters as well.
B.3 Few shot transfer learning for clean data
Fig. B.3.1: Violin plot, few-shot learning, clean data, 301 data points: of the predicted value of the unknown parameters. 301 data points are considered for each of the known parameters (for each one minute segment) and no noise is considered here. The violin plots are clipped at the maximum and minimum value of the predicted parameters. The dots over the violin plots show the swarm plot indicate predicted value for individual runs. The grey region shown between the dotted lines indicates the $\pm$ 7.5% error band. The middle dotted line indicates the true value of the unknown parameter, the top dotted line indicates 1.075 times the true value while the bottom dotted line indicates 0.925 times the true value.
Fig. B.3.2: Predicted states, Few-shot learning, Clean data, 151 data points, ensemble: Ensemble prediction for state variable and $x_{r}$ and $T_{1}$ for clean data with 151 data points in the data loss predicted using few-shot transfer learning. The ensembles are calculated over 30 independent runs for each time segment.
Fig. B.3.3: Predicted states, Few-shot learning, Clean data, 151 data points, ensemble: Ensemble prediction for different variables for clean data with 151 data points in the data loss predicted using few-shot transfer learning. The ensembles are calculated over 30 independent runs for each time segment.
Fig. B.3.4: Predicted states, Few-shot learning, Clean data, 151 data points, ensemble: Ensemble prediction for the empirical formulas for clean data with 151 data points in the data loss predicted using few-shot transfer learning. The ensembles are calculated over 30 independent runs for each time segment.
B.4 Few shot transfer learning for noisy data
Fig. B.4.1: Predicted states, Few-shot learning, Noisy data, 301 data points, ensemble: Ensemble prediction for state variable and $x_{r}$ and $T_{1}$ for noisy data with 301 data points in the data loss predicted using few-shot transfer learning. The ensembles are calculated over 30 independent runs for each time segment.
Fig. B.4.2: Predicted states, Few-shot learning, Noisy data, 301 data points, ensemble: Ensemble prediction for different variables for noisy data with 301 data points in the data loss predicted using few-shot transfer learning. The ensembles are calculated over 30 independent runs for each time segment.
Fig. B.4.3: Predicted states, Few-shot learning, Noisy data, 301 data points, ensemble: Ensemble prediction for the empirical formulas for noisy data with 301 data points in the data loss predicted using few-shot transfer learning. The ensembles are calculated over 30 independent runs for each time segment.
Appendix C Additional Tables
Table C.1: Input output structure of DeepONets for EGR ($\tilde{u}_{egr1}$, $\tilde{u}_{egr1}$) and VGT ($\tilde{u}_{vgt}$) states. As shown in Fig. 5, the DeepONet ($\mathcal{D}_{1}$) for EGR states has four branches and two outputs. The DeepONet ($\mathcal{D}_{1}$) for VGT state has two branches and one output. The input to the trunk is $t=0-60$ sec and scaled between $[-1,1]$. A schematic diagram of the DeepONet is shown in Fig. 5.
Trunk Branch-1 Branch-2 Branch-3 Branch-4 Output
For EGR states $\tilde{u}_{egr1}$ and $\tilde{u}_{egr2}$
$\begin{matrix}t=t_{0}=0.00\\ t=t_{1}=0.02\\ \cdot\\ \cdot\\ t=t_{n}=60.0\end{matrix}$ $\begin{matrix}u_{degr,0}&0&0&\dots&0\\ u_{degr,1}&u_{degr,0}&0&\dots&0\\ \cdot&\cdot&\cdot&\cdots&\cdot\\ \cdot&\cdot&\cdot&\cdots&\cdot\\ u_{degr,n}&u_{degr,n-1}&u_{degr,n-2}&\dots&u_{degr,0}\\ \end{matrix}$ $\begin{matrix}\tilde{u}_{egr1,0}\\ \tilde{u}_{egr1,0}\\ \cdot\\ \cdot\\ \tilde{u}_{egr1,0}\end{matrix}$ $\begin{matrix}u_{degr,0}&0&0&\dots&0\\ u_{degr,1}&u_{degr,0}&0&\dots&0\\ \cdot&\cdot&\cdot&\cdots&\cdot\\ \cdot&\cdot&\cdot&\cdots&\cdot\\ u_{degr,n}&u_{degr,n-1}&u_{degr,n-2}&\dots&u_{degr,0}\\ \end{matrix}$ $\begin{matrix}\tilde{u}_{egr2,0}\\ \tilde{u}_{egr2,0}\\ \cdot\\ \cdot\\ \tilde{u}_{egr2,0}\end{matrix}$ $\begin{matrix}\tilde{u}_{egr1,0}\\ \tilde{u}_{egr1,1}\\ \cdot\\ \cdot\\ \tilde{u}_{egr1,n}\end{matrix}$ $\begin{matrix}\tilde{u}_{egr2,0}\\ \tilde{u}_{egr2,1}\\ \cdot\\ \cdot\\ \tilde{u}_{egr2,n}\end{matrix}$
For VGT states $\tilde{u}_{vgt}$
$\begin{matrix}t=t_{0}=0.00\\ t=t_{1}=0.02\\ \cdot\\ \cdot\\ t=t_{n}=60.0\end{matrix}$ $\begin{matrix}u_{dvgt,0}&0&0&\dots&0\\ u_{dvgt,1}&u_{dvgt,0}&0&\dots&0\\ \cdot&\cdot&\cdot&\cdots&\cdot\\ \cdot&\cdot&\cdot&\cdots&\cdot\\ u_{dvgt,n}&u_{dvgt,n-1}&u_{dvgt,n-2}&\dots&u_{dvgt,0}\\ \end{matrix}$ $\begin{matrix}\tilde{u}_{vgt,0}\\ \tilde{u}_{vgt,0}\\ \cdot\\ \cdot\\ \tilde{u}_{vgt,0}\end{matrix}$ NA NA $\begin{matrix}\tilde{u}_{vgt,0}\\ \tilde{u}_{vgt,1}\\ \cdot\\ \cdot\\ \tilde{u}_{vgt,n}\end{matrix}$